### v2

# Un-mixing Test-time Adaptation under Heterogeneous Data Streams

Abstract
Deploying deep models in real-world scenarios remains challenging due to significant performance drops under distribution shifts between training and deployment environments. Test-Time Adaptation (TTA) has recently emerged as a promising solution, enabling on-the-fly model adaptation without access to source data. However, its effectiveness degrades significantly in the presence of complex, mixed distribution shifts – common in practical settings – where multiple latent domains coexist. Adapting under such intrinsic heterogeneity, especially in unlabeled and online conditions, remains an open and underexplored challenge. In this paper, we study TTA under mixed distribution shifts and move beyond conventional homogeneous adaptation paradigms. By revisiting TTA from a frequency-domain perspective, we observe that distribution heterogeneity often manifests in Fourier space – for instance, high-frequency components tend to carry domain-specific variations. This motivates us to perform domain-aware separation using high-frequency texture cues, making diverse shift patterns more tractable. To this end, we propose FreDA, a novel Frequency-based Decentralized Adaptation framework that decomposes globally heterogeneous data into locally homogeneous components in the frequency domain. It further employs decentralized learning and augmentation strategies to robustly adapt under complex, evolving shifts. Extensive experiments across various environments (corrupted, natural, and medical) demonstrate the superiority of our proposed framework over the state-of-the-arts.
Index Terms:
Test-time Adaptation, Transfer Learning
I Introduction
Deep learning models often suffer significant performance degradation when deployed in environments where the data distribution differs from that of the training set – a challenge known as domain shift [1, 2]. Recently, Test-Time Adaptation (TTA) [3, 4, 5, 6, 7, 8, 9, 10] has emerged as a promising solution by refining model parameters to better align with the encountered data at inference time. It leverages the incoming data stream for real-time adjustments without the need for retraining on a labeled dataset, enabling swift model adaptation to unpredictable data characteristics during deployment.
Despite their success, current TTA models often are limited to ideal testing conditions, typically involving homogeneous test samples with similar types of distribution shifts. In reality, distribution shifts are mixed, overlapping, and even conflicting [11, 12, 13, 14, 15]. For instance, photo management software handles diverse corruptions, such as noise, blur, compression; medical imaging platforms manage inconsistencies from varied acquisition methods; and autonomous driving systems face fluctuating conditions like lighting, weather, and road types.
Such complexity poses serious challenges for existing methods. While recent efforts have extended TTA to continually changing environments [5], they typically assume a uniform target domain at each time step, as illustrated in the continual setting in Figure 1 (a). Some methods periodically reset the model to its source pre-trained state [7, 9] to mitigate the cumulative effect of sequential distribution shifts on adaptation. Others down-weight outlier target samples [6, 10] to suppress the impact of abrupt or rare shifts. Although these strategies aim to improve robustness under dynamic conditions, they fall short when facing the entangled / mixed real-world distributions (see Figure 1 (b)). Their shortcomings stem from a shared limitation: they lack the capacity to disentangle or localize domain variations. As a result, they often overfit to dominant patterns or suffer from catastrophic forgetting, failing to be deployed in realistic, heterogeneous data streams.
Figure 1: (a) Illustration of three TTA scenarios with single, continual, and mixed domain shifts. (b) Classification error rate on CIFAR-10-C across different TTA settings: performance of conventional models drops sharply under mixed domain shifts.
Capturing such intrinsic heterogeneity under unlabeled, online conditions is particularly challenging, and effective adaptation in this setting remains an open problem. To tackle this, we propose shifting from coarse, homogeneous adaptation to a fine-grained strategy that explicitly disentangles heterogeneous shifts. Our approach leverages the frequency information to disentangle domain variations, which naturally separates data features across frequency bands: high-frequency components capture fine details like edges and textures, while low-frequency components represent global structures such as shape and illumination (Section IV). By decomposing inputs accordingly, we better characterize distributional variations and identify diverse shifts. Moreover, the Fourier transform operates directly on raw pixel-level inputs, which shields the method from performance degradation under large domain gaps, and also enhancing adaptability and deployment efficiency in real-world applications.
Building upon this insight, we introduce a framework, termed Frequency-based Decentralized Adaptation (FreDA). FreDA begins by partitioning incoming data in the Fourier domain, where high-frequency components are used to group samples with similar shift characteristics. This transforms globally heterogeneous inputs into locally homogeneous subsets before any adaptation takes place. Based on these partitions, FreDA deploys multiple local models, each assigned to a specific subset. These models adapt independently to their respective data streams while periodically synchronizing through parameter exchange. The aggregated parameters form a shared base model, which is then used to reinitialize the local models, ensuring continual and coordinated adaptation across domains. This approach not only alleviates clustering errors via collaborative knowledge sharing but also allows local models to capture diverse distribution shifts. To further enhance robustness, we introduce a novel Fourier-based augmentation scheme that improves sample quality and strengthens adaptation to shift-specific characteristics.
To summarize, the main contributions of this work are three-fold:
  • We identify a key limitation of most existing TTA methods – their neglect of real-world data heterogeneity – which leads to suboptimal performance when confronted with mixed and diverse distribution shifts.
  • We propose FreDA, a frequency-based decentralized adaptation framework that leverages spectral decomposition and localized adaptation to effectively address heterogeneous distribution shifts at test-time.
  • We validate FreDA through extensive experiments on corrupted, natural, and medical benchmarks, demonstrating consistent improvements over state-of-the-art methods across various TTA scenarios.
II Connections to Previous Studies
II-A Transfer Learning, Domain Adaptation, Test-time Training, and Test-time Adaptation
Deep neural networks often experience performance degradation when deployed in environments different from their training settings, due to distribution shifts between source (training) and target (testing) domains. This issue is extensively studied under the umbrella of transfer learning [16, 17], which aims to transfer knowledge across domains. A key branch of transfer learning is domain adaptation (DA) [18, 19, 20, 21], where models trained on a labeled source domain are adapted to a target domain. Depending on the availability of source and target data, DA can be classified into: Supervised DA using labeled data in both source and target domains during training; Unsupervised DA, which relies on labeled source data and unlabeled target data; and Source-free DA, that adapts to the target domain without access to source data, typically due to privacy or transmission constraints.
TABLE I: Comparison of different transfer learning settings.
Topic Source Data Target Labels Online Adaptation Model Agnostic
Supervised Domain Adaptation $\checkmark$ $\checkmark$ $\times$
Unsupervised Domain Adaptation $\checkmark$ $\times$ $\times$
Source-free Domain Adaptation $\times$ $\times$ $\times$
Test-time Training $\times$ $\times$ $\checkmark$ $\times$
Test-time Adaptation $\times$ $\times$ $\checkmark$ $\checkmark$
Note: TTA is the most practical setting for real-world deep model deployment.
While traditional DA methods assume access to the whole target distribution during adaptation, such access is not always possible in real-world deployment. To address this, two prominent test-time adaptation paradigms have been proposed: Test-time Training (TTT) [22, 23, 24] introduces an auxiliary task – often self-supervised – during pre-training. At test time, the model optimizes this auxiliary objective to adapt to the target distribution accordingly; Test-time Adaptation (TTA) [3, 7, 8] poses a more demanding and practical scenario, where the model must adapt on-the-fly to the test stream without any prior modification during training. This setting emphasizes the need for rapid, real-time model updates to effectively capture the continuously incoming data. A brief comparison of these adaptation settings is summarized in TABLE I. Our work falls under the TTA setting, where we propose a generalizable and robust adaptation framework that maintains strong performance across diverse distribution shifts and corruptions – without relying on source data or altering the pre-training process.
II-B Non-i.i.d. Test-time Adaptation
While conventional TTA methods primarily focus on single domain shifts under the idealized independent and identically distributed (i.i.d.) assumption, real-world deployment often violates these conditions due to inherent data heterogeneity. This discrepancy has motivated recent efforts to extend TTA to more realistic settings, with two primary challenges:
II-B1 Mixed Domains
Modern TTA methods designed for continual domain shifts [5, 25, 6, 9, 10] have pushed beyond the single adaptation setting. However, these approaches assume that the target domain is uniform at each time point – a special case of mixed domains where the change occurs across distinct time periods. In reality, data streams often involve genuinely mixed domains with heterogeneous distributions present simultaneously. Existing approaches primarily focus on stabilizing model updates through periodic parameter resets [7, 9] or importance weighting of target samples [6, 9, 10], but these techniques are ill-equipped to disentangle and manage the complex interleaving of domain shifts, leading to degraded performance. In contrast, our work adopts a data-centric perspective by explicitly addressing mixed domains via frequency-space decomposition, which facilitates domain separation and proactive distribution alignment prior to model adaptation. While recent efforts like [7] acknowledge mixed distribution challenges under the broader “Dynamic Wild World” paradigm, their unified treatment of diverse real-world factors does not systematically address data heterogeneity. Conversely, our study directly targets the core issue of mixed domains in TTA, introducing tailored strategies for disentangling and adapting to co-occurring heterogeneous distributions.
Figure 2: (a) Visual heterogeneity across five healthcare centers in Camelyon17 dataset: Example patches showcase diverse domain shifts that characterize mixed target domains. (b) Conventional sample latent features from pretrained models fail to separate different target subdomains, showing significant overlap. (c) High-frequency information enable distinct separation of target subdomains, contrasting with the conventional ones.
II-B2 Dependent Sampling
The second challenge stems from class-level dependencies introduced by temporal data sampling. This topic has garnered significant research attention, with a growing body of work [25, 26, 27, 28, 29, 30] actively addressing it through strategies such as pseudo-label-based rebalancing, or extended observation windows. These approaches aim to correct the temporal imbalance caused by skewed class distributions over time. However, unlike this these studies that focus on class-level imbalance, our work diverges by targeting sample-level heterogeneity, which remains underexplored. Specifically, we address style diversity and distributional variation at test-time – a challenge that persist even under class-balanced conditions – to improve model robustness under mixed-domain scenarios.
III Problem Definition of mixed domain shifts
Test-time adaptation (TTA) aims to adjust a model $q_{\mathbf{\theta}}(y|x)$, initially trained on a source dataset $\mathcal{D}_{s}=\{(x,y)\sim p_{s}(x,y)\}$, to a target domain $\mathcal{D}_{t}=\{(x,y)\sim p_{t}(x,y)\}$ on a data stream without accessing source data or target labels. TTA handles covariate shift by assuming $p_{s}(y|x)=p_{t}(y|x)$ while $p_{s}(x)\neq p_{t}(x)$. This challenge intensifies when $\mathcal{D}_{t}$ contains multiple non-i.i.d sub-distributions $p_{t_{i}}(x)$, such as
$p_{t}(x)\leftarrow\{p_{t_{1}}(x),p_{t_{2}}(x),\dots,p_{t_{N}}(x)\}.$
Specifically, we have $\mathcal{D}_{t}=\mathcal{D}_{t_{1}}\cup\mathcal{D}_{t_{2}}\cup\cdots\cup\mathcal{D}_{t_{N}}$ where $x\in\mathcal{D}_{t_{1}}$ satisfying $x\sim p_{t_{1}}(x)$. This scenario requires the model $q_{\mathbf{\theta}}(y|x)$ to effectively handle the heterogeneous and evolving target distribution to maintain robust performance. TTA strategies must therefore refine the model to optimize its predictive accuracy across these diverse sub-domains, ensuring consistent and reliable performance amidst significant distributional variability.
Methods C10 C100 IN
TBN (Centralized) 33.8 45.8 82.5
TBN (Decentralized) 28.5 43.2 77.6
(a) Centralized Adaptation
(b) Decentralized Adaptation
Figure 3: Experimental Result (error rate) and t-SNE feature visualization using TBN [31] as the pseudo-labeling method under mixed distribution shifts, comparing centralized and decentralized adaptation. Different colors represent different classes. (a) Centralized Adaptation: Global BN parameters are applied to the entire batch. (b) Decentralized Adaptation: Localized BN parameters are tailored for clusters separated based on high-frequency features. C10, C100, IN respectively denotes CIFAR-10-C, CIFAR-100-C and ImageNet-C datasets.
IV TTA under Mixed Distribution Shifts: A Fourier Perspective
While Test-Time Adaptation (TTA) methods excel under single-type distribution shifts, their performance degrades catastrophically when facing mixed distribution shifts – a ubiquitous challenge in real-world deployment. As demonstrated on CIFAR-10-C (Figure 1 (b)), leading methods like TENT [3], SAR [7] and RoTTA [25] suffer 16% average accuracy drop compared to single-shift scenarios. Conventional TTA paradigms attempt to mitigate this through passive robustness measures, such as entropy regularization [3], gradient masking [7], or model resetting [9]. These methods treat mixed distribution shifts as irreducible noise, deferring corrective actions until model failures occur – a reactive paradigm that inevitably compromises adaptation efficacy.
To address this, we propose a paradigm shift to proactive distribution management – instead of demanding models to handle arbitrary mixtures, we first resolve the heterogeneity via frequency-space decomposition. Our key insight is that mixed distributions exhibit spectral signatures that can be disentangled before model adaptation begins. As evidenced in Figure 2 (c), high-frequency components naturally separate distinct shifts without relying on error-prone model features (see Figure 2 (b). This theoretically grounded decomposition converts the ill-posed mixed-shift adaptation into multiple well-conditioned homogeneous-shift problems, effectively create an ideal adaptation environment from chaotic real-world streams. Crucially, this decomposition enables a decentralized adaptation where we can deploy domain-specific models tailored to high-frequency-separated clusters. Take TBN [31] as an example: while its centralized variant recalculates batch normalization (BN) statistics globally (treating heterogeneous data as a single distribution), our decentralized adaptation computes localized BN parameters within each spectrally disentangled cluster. As shown in Figu 3, this context-aware estimation preserves semantic discriminability: samples exhibit tighter within-class cohesion and sharper between-class boundaries compared to the blurred separability of centralized adaptation. Building on these insights, the following section details how leveraging the frequency domain enhances TTA methods for realistic scenarios with mixed distribution shifts.
Figure 4: Illustration of Frequency-based Decentralized Learning.
V Frequency-based Decentralized Adaptation
The previous discussion highlights how heterogeneity within target distributions can hinder effective model adaptation. This naturally raises the question: How can we manage such distributional heterogeneity to enable better adaptation? As outlined earlier, distinguishing samples associated with different distribution shifts offers a promising pathway. Building on this insight, we address the TTA problem by leveraging high-frequency components in the data and propose a novel framework termed Frequency-based Decentralized Adaptation (FreDA). It partitions target samples into multiple homogeneous subdomains in the Fourier space, facilitating more accurate model adaptation. This is further enhanced by a frequency-based augmentation strategy that enriches each subdomain with diverse samples, strengthening model robustness.
V-A Frequency-based Decentralized Learning
Fourier transform offers an effective method to extract different frequency components from images, with high-frequency information particularly useful for capturing fine-grained details such as texture and noise. These details often highlight subtle variations among different distribution shifts. Based on this insight, we propose a module termed Frequency-based Decentralized Learning. It leverages frequency information extracted from the pixel space to systematically partition data into multiple homogeneous subsets, enabling multiple local models to specialize in capturing each distribution shift individually. Concurrently, our method enables collaborative learning by allowing periodic parameter sharing among these local models via an enhanced base model, thereby boosting overall adaptability to diverse distribution shifts.
V-A1 Frequency Feature Extraction
We start by extracting frequency domain features from the input images. Let $\mathbf{X}\in\mathbb{R}^{n\times c\times h\times w}$ denote a batch of input images, where $n$ is the batch size, $c$ is the number of channels, $h$ and $w$ are the height and width. We first apply a Fourier transform $\mathcal{F}$ to each image $\mathbf{x}_{i}$. This transform converts the image from the spatial domain to the frequency domain, producing a complex-valued representation $\mathcal{F}(\mathbf{x}_{i})\in\mathbb{C}^{h\times w\times c}$ that contains both real $R(\mathbf{x}_{i})$ and imaginary part $I(\mathbf{x}_{i})$. Next, we compute the amplitude spectrum $A(\mathbf{x}_{i})(u,v)$ using $A(x)(u,v)=\sqrt{R^{2}(x)(u,v)+I^{2}(x)(u,v)}$. It reveals the intensity of the frequency content, e.g., high-frequency amplitudes highlight edges and fine details while low-frequency amplitudes emphasize the overall structure and gradual changes. Then, we filter out low-frequency elements using mask $M(u,v)=\mathbbm{1}\left(\left(u<\frac{h}{4}\lor u>\frac{3h}{4}\right)\lor\left(v<\frac{w}{4}\lor v>\frac{3w}{4}\right)\right)$ to emphasize the high-frequency components $G(x)(u,v)$ that are more likely to indicate shifts in distribution:
$G(x)(u,v)=A(x)(u,v)\cdot M(u,v).$ (1)
V-A2 Frequency-Based Clustering
We then employ a clustering algorithm (e.g., K-means) to partition the frequency features into $K$ clusters, each corresponding to a different type of distribution shift. The process is formalized as:
$\min_{\mathbf{C},\mathbf{Z}}\sum_{i=1}^{n}\left\|\mathbf{A}_{hf,i}-\mathbf{C}_{\mathbf{Z}_{i}}\right\|^{2}_{2},$ (2)
where $\mathbf{A}_{hf,i}=\text{vec}(G(x_{i}))$, $\mathbf{C}\in\mathbb{C}^{K\times d}$, $\mathbf{Z}\in\{1,\dots,K\}^{n}$ denotes the 1D high-frequency component of the amplitude spectrum, the centroids of the clusters and the cluster assignments for each image. $hf$ refers to high-frequency components, and $d=h\times w\times c$ is flattened dimension.
V-A3 Decentralized Fine-tuning
Test-time fine-tuning is then decentralized across these clusters, allowing for specialized adaptation within each subgroup: For each cluster $k$, we adapt a specialized model $q_{\theta_{k}}(y|x)$ that is fine-tuned using only the data within that cluster:
$\theta_{k}^{*}=\arg\min_{\theta_{k}}\mathbb{E}_{x\sim{p}_{t,k}}\left[\mathcal{L}(q_{\theta_{k}}(x))\right],$ (3)
where ${p}_{t,k}$ represents the data distribution within cluster $k$, and $\mathcal{L}$ is the loss function. The predictions of each iteration are collected and sorted after the local fine-tuning.
V-A4 Base Model Enhancement
To integrate knowledge from all subnetworks and prevent degradation on specific subdomains, we periodically aggregation their parameters at intervals of time $T$:
$\mathbf{\theta}_{\text{base}}=\sum_{k=1}^{K}\left(\frac{|\mathcal{D}_{k}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}|}\mathbf{\theta}_{k}\right),$ (4)
where $|\mathcal{D}_{k}|$ denotes sample number in cluster $k$. This aggregation step combines the parameter updates from each subnetwork proportionally to its cluster size. The updated parameters $\mathbf{\theta}_{\text{base}}$ are then distributed back to each subnetwork, initializing them for the next batch of training: $\mathbf{\theta}_{k}\leftarrow\mathbf{\theta}_{\text{base}}$.
V-B Frequency-based Augmentation
While decentralized learning effectively mitigates batch-level heterogeneity, it may falls short in accurately characterizing individual distribution shifts – largely due to limited batch size and noise introduced by coarse clustering. To improve target data quality, TTA methods commonly adopt data augmentation as a practical strategy to enhance model generalization. However, conventional augmentation techniques – typically borrowed from standard computer vision tasks such as rotation, cropping, and mixup – are primarily tailored for single-shift scenarios. They often fail to provide the targeted, distribution shifts-aware augmentation as required under mixed domain shifts. To this end, we propose a frequency-based augmentation strategy tailored for TTA in mixed-shift environments. Specifically, our method perturbs the amplitude components of target samples in the Fourier space, enabling targeted augmentation with respect to each sample’s underlying distribution. By manipulating the frequency-domain characteristics, our approach generates diverse yet distribution-consistent variations, thereby enriching subdomain representations and facilitating robust model adaptation
Algorithm 1 Framework of Frequency-based Decentralized Learning and Augmentation
1:Step $t$, Input batch $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}\in\mathbb{R}^{n\times h\times w\times c}$, Pretrained source model $q_{\theta}$, Initialize Feature Repository and Local Sample Pool $\mathcal{R},\mathcal{S}_{k}\leftarrow\emptyset$, CLUSTER_NUM $K$, KMEANS_SIZE $N$, COMM_INTERVAL $f$;
2:Step 1: Extract Frequency Features
3:for $i=1$ to $n$ do
4:  $\mathbf{A}_{hf,i}\leftarrow\text{vec}(\text{G}(\mathbf{x}_{i}))$triangleright Extract high-freq components
5:end for
6:Step 2: Dynamic Clustering
[…]
10:Step 3: Local Model Training
11:for cluster $k\in\{1,\ldots,K\}$ do
12:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}\cup\{x_{i}\mid Z_{i}=k\}$triangleright Gather samples for cluster $k$
13:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}[(|\mathcal{S}_{k}|-n+1):]$triangleright Keep the last $batch\_size=n$ entries
14:  $\mathcal{S}_{k}^{\prime}\leftarrow\text{select\_samples}(\mathcal{S}_{k})$triangleright Select samples (Eq. 5)
15:  for each $x_{i}\in\mathcal{S}_{k}^{\prime}$ do
16:   $\tilde{x}_{i}\leftarrow\text{augment}(x_{i})$triangleright Augment data (Eq. 7)
17:   $\text{Train}(q_{\theta_{k}},x_{i},\tilde{x}_{i})$triangleright Train local model (Eq. 3)
18:  end for
19:end for
20:Step 4: Compile Predictions
[…]
22:Step 5: Base Model Enhancement
23:If t % $f$ == 0 : $\triangleright$ Model Communication with interval $f$ (Eq.4)
[…]
V-B1 Sample Selection Mechanism
We first select the reliable samples in each local model leveraging a criterion derived from the weighted entropy framework used in ETA [6] based on two primary conditions:
$\text{Cri}=\mathbbm{1}\left[\left(H(\mathbf{y}_{t})<H_{0}\right)\wedge\left(|\text{cos}(\mathbf{y}_{t},\bar{\mathbf{y}}_{t-1})|<\epsilon\right)\right].$ (5)
The entropy $H(\mathbf{y}_{t})$ measures the uncertainty in the current predictions. The cosine similarity $\text{cos}(\mathbf{y}_{t},\bar{\mathbf{y}}_{t-1})$ denotes the deviation between the current sample’s class probabilities $\mathbf{y}_{t}$ and the aggregated class probabilities $\bar{\mathbf{y}}_{t-1}$. $\epsilon$ is the threshold for cosine similarity, and $H_{0}$ is the fixed entropy threshold. This ensures that selected samples exhibit significant deviations from previous predictions in class distribution and lower prediction uncertainty.
V-B2 Frequency-Based Augmentation
The augmentation process involves perturbing the amplitude spectrum. Let $A(x_{i})$ represent the amplitude spectrum of a selected sample $x_{i}$. To generate a perturbed amplitude spectrum $\tilde{A}(x_{i})$, we apply a random Gaussian perturbation:
$\tilde{A}(x_{i})=\left(1+\alpha\cdot\Delta\right)\cdot A(x_{i}),$ (6)
where $\Delta\sim\mathcal{N}(0,\sigma^{2})$ is a perturbation matrix sampled from a Gaussian distribution, and $\alpha$ is a scaling factor. Then, the augmented sample $\tilde{x}_{i}$ is reconstructed via the inverse Fourier transform $\mathcal{F}^{-1}$ to the perturbed amplitude spectrum, combined with the original phase spectrum $P(x_{i})$:
$\tilde{\mathbf{X}}_{i}=\mathcal{F}^{-1}\left(\tilde{A}(x_{i}),P(x_{i})\right).$ (7)
V-B3 Loss Function
The training objective combines the entropy loss of the selected samples with a consistency loss from the augmented samples. The total loss is defined as:
$\mathcal{L}_{\text{total}}=\frac{1}{n}\sum_{i=1}^{n}H(\mathbf{y}_{i})+\lambda\cdot\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{\text{con}}\left(\hat{\mathbf{y}}_{i},\tilde{\mathbf{y}}_{i}\right),$ (8)
where the entropy loss $H(\mathbf{y}_{i})$ for the original sample $x_{i}$ is given by $H(\mathbf{y}_{i})=-\sum_{j=1}^{C}\mathbf{y}_{i,j}\log\mathbf{y}_{i,j}$ with $\mathbf{y}_{i}$ being the predicted probability over the $C$ classes, and the consistency loss $\mathcal{L}_{\text{con}}\left(\hat{\mathbf{y}}_{i},\tilde{\mathbf{y}}_{i}\right)=-\sum_{j=1}^{C}\hat{\mathbf{y}}_{i,j}\log\tilde{\mathbf{y}}_{i,j}$ is defined as the cross-entropy between the prediction $\tilde{\mathbf{y}}_{i}$ of the augmented sample $\tilde{x}_{i}$ and the pseudo-label $\hat{\mathbf{y}}_{i}$ from the original sample.
V-C Overall Framework
We provide the overall pipeline algorithm of FreDA in Algorithm 1. During implementation, we leverage a memory bank strategy [25, 4, 32, 33] that is updated in real time. This design serves two purposes: 1) to ensure accurate clustering—since an overly small batch could impede the effective separation of data – and 2) to maintain the number of samples processed by the local model consistent with the original batch size, thereby preventing performance degradation due to a drastic reduction in batch size (e.g., reducing by a factor of the cluster number).
VI Theoretical Insights
To facilitate theoretical understanding, we leverage the expansion-based analysis framework proposed in [34]. We first revisit the core concepts and then highlight the theoretical advantages of our method under mixed domain shifts.
VI-A Expansion Theory
Definition 1 ($(a,c)$-expansion).
[…]
$\mathcal{P}_{i}(\mathcal{N}(S))\geq\min(c\mathcal{P}_{i}(S),1)$
where $\mathcal{N}(S)$ refers to the neighborhood of $S$ under data augmentations.
Definition 2 (Separation).
$P$ is $(\mu,r)$-separated if:
$\mathbb{E}_{x\sim\mathcal{P}}\left[\max_{x^{\prime}\in B_{r}(x)}\mathbf{1}(G(x)\neq G(x^{\prime}))\right]\leq\mu$
where $B_{r}(x)$ is an $\ell_{2}$-ball of radius $r$.
The core theorem from [34] states:
Theorem 1 (Pseudo-label Denoising).
Under $(a,c)$-expansion and $(\mu,r)$-separation, any classifier $G$ of:
$\min_{G}\frac{c+1}{c-1}\mathcal{L}_{pl}(G)+\frac{2c}{c-1}\mathcal{R}_{B}(G)$
achieves error:
$\mathrm{Err}(G)\leq\frac{2}{c-1}\mathrm{Err}(G_{pl})+\frac{2c}{c-1}\mu$
where $\mathcal{L}_{pl}$ is pseudo-label loss and $\mathcal{R}_{B}$ is consistency regularizer.
VI-B Analysis of Our Method
Our method reduces both terms in Theorem 1 to achieve a tighter bound through two mechanisms:
VI-B1 Frequency-Coherent Partitioning.
Let $\{P_{t_{k}}\}_{k=1}^{K}$ be the $K$ sub-domains identified by frequency-based partitioning. For measurable sets $\{S_{k}\}_{k=1}^{K}$ where $S_{k}\subseteq X_{k}$ (the support of $P_{t_{k}}$), define:
$\mathrm{Err}(G_{pl}^{k})\leq\mathrm{Err}(G_{pl})-\Delta_{k}\quad$
where $\Delta_{k}=\mathcal{P}_{t_{k}}\left(E_{c}\right)$ and $E_{c}=\{x\in X_{k}\mid G_{pl}(x)\neq y(x)\}$. This contributes to the reduction of the first term in Theorem 1, which is further supported by experimental results in Figure 3.
VI-B2 Augmentation-Induced Expansion
Our frequency augmentation expands neighborhoods:
$\widehat{\mathcal{N}}(S)=\mathcal{N}_{\text{base}}(S)\cup\left\{x^{\prime}:\inf_{x\in S}|A(\mathcal{F}(x^{\prime}))-A(\mathcal{F}(x))|_{2}\leq\epsilon\right\}$
where $\mathcal{N}_{\text{base}}(S)$ is the base augmentation neighborhood. It is worth noting that, unlike most methods, we do not apply base augmentation. The term $\mathcal{N}_{\text{base}}(S)$ is retained in the formulation for theoretical completeness. Our choice of this representation is intended to highlight the expansion gain brought by incorporating frequency augmentation.
The expansion gain becomes:
$\begin{split}\widehat{c}&=c_{\text{base}}+\gamma,\\ \gamma=\inf_{S:P(S)\leq a}&\frac{P(\widehat{\mathcal{N}}(S)\setminus\mathcal{N}_{\text{base}}(S))}{P(S)}\end{split}$
Substituting into Theorem 1, the error becomes:
$\begin{split}\mathrm{Err}(G)&\leq\underbrace{\frac{2}{(c_{\text{base}}+\gamma)-1}\left(\mathrm{Err}(G_{pl})-\sum_{k=1}^{K}\Delta_{k}\right)}_{\text{Reduced by both terms}}\\ &+\underbrace{\frac{2(c_{\text{base}}+\gamma)}{(c_{\text{base}}+\gamma)-1}\mu}_{\text{Reduced by }O(\gamma/c^{2})}\end{split}$
Notation Summary
  • $c$: Expansion factor (Def. 1).
    Class-wise connectivity metric: Larger $c$ implies stronger neighborhood propagation of local consistency to global predictions.
  • $\mu$: Inter-class separation probability (Def. 2).
    Robustness measure: Probability of different classes having overlapping neighborhoods under perturbations, lower $\mu$ indicates clearer decision boundaries.
  • $\mathcal{R}_{B}$: Consistency regularizer (Thm. 1).
    Stability term: Penalizes prediction inconsistency between original inputs and their augmented variants.
  • $\Delta_{k}$: Error reduction in $k$-th sub-domain.
    Specialization gain: Reduced pseudo-label noise in sub-domain $k$ due to frequency-based partitioning.
  • $\gamma$: Expansion gain from frequency augmentation.
    Neighborhood enhancement: Additional expansion capability measured as the relative increase of augmented neighborhoods.
TABLE II: Classification error rate ($\downarrow$) on CIFAR-10-C, CIFAR-100-C, and ImageNet-C (IN-C) under Mixed Distribution Shifts.
Baseline & Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
CIFAR-10-C (WRN-28) 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.4 30.3 43.5
TBN 45.5 42.8 59.7 34.2 44.3 29.8 32.0 19.8 21.1 21.5 9.3 27.9 33.1 55.5 30.8 33.8
TENT (ICLR 21’) 73.5 70.1 81.4 31.6 60.3 29.6 28.5 30.8 35.3 25.7 13.6 44.2 32.6 70.2 34.9 44.1
ETA (ICML 22’) 36.2 33.3 52.3 22.9 38.9 22.4 20.5 19.5 19.7 20.4 11.3 35.4 26.6 38.8 25.1 28.2
AdaContrast (CVPR 22’) 36.7 34.3 48.8 18.2 39.1 21.1 17.7 18.6 18.3 16.8 9.0 17.4 27.7 44.8 24.9 26.2
CoTTA (CVPR 22’) 38.7 36.0 56.1 36.0 36.8 32.3 31.0 19.9 17.6 27.2 11.7 52.6 30.5 35.8 25.7 32.5
SAR (ICLR 23’) 45.5 42.7 59.6 34.1 44.3 29.7 31.9 19.8 21.1 21.5 9.3 27.8 33.0 55.4 30.8 33.8
RoTTA (CVPR 23’) 60.0 55.5 70.0 23.8 44.1 20.7 21.3 20.2 22.7 16.0 9.4 22.7 27.0 58.6 29.2 33.4
RDumb (NeurIPS 23’) 34.9 32.3 49.4 23.3 38.2 23.3 20.7 19.9 19.3 20.7 11.2 29.3 26.7 41.5 25.2 27.7
DeYO (ICLR 24’) 45.8 42.3 65.7 21.3 41.8 25.1 19.5 21.1 19.6 19.2 12.3 21.8 28.5 39.3 28.0 30.1
UnMix-TNS (ICLR 24’) 50.0 44.4 44.3 34.4 48.2 32.7 30.0 35.5 35.9 47.5 28.1 38.7 43.9 40.0 43.3 39.8
FreDA (ours) 23.1 22.2 32.2 18.7 41.6 18.8 16.8 17.9 19.9 16.9 9.8 13.2 29.1 35.4 28.6 22.9
CIFAR-100-C (ResNeXt-29) 73.0 68.0 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
TBN 62.7 60.7 43.1 35.5 50.3 35.7 34.4 39.9 51.5 27.5 45.5 42.3 72.8 46.4 45.8 45.8
TENT (ICLR 21’) 95.6 95.2 89.2 72.8 82.9 74.4 72.3 78.0 79.7 84.7 71.0 88.5 77.8 96.8 78.7 82.5
ETA (ICML 22’) 42.6 40.3 34.1 30.3 42.4 32.0 29.4 35.6 35.8 44.1 30.2 41.8 36.9 38.9 40.9 37.0
AdaContrast (CVPR 22’) 54.5 51.5 37.6 30.7 45.4 32.1 30.3 36.9 36.5 45.3 28.0 42.7 38.2 75.4 41.7 41.8
CoTTA (CVPR 22’) 54.4 52.7 49.8 36.0 45.8 36.7 33.9 38.9 35.8 52.0 30.4 60.9 40.2 38.0 41.1 43.1
SAR (ICLR 23’) 75.8 72.7 41.1 29.2 45.2 31.1 28.9 36.7 37.7 43.9 29.3 41.8 37.1 89.2 42.4 45.5
RoTTA (CVPR 23’) 65.0 62.3 39.3 33.4 50.0 34.2 32.6 36.6 36.5 45.0 26.4 41.6 40.6 89.5 48.5 45.4
RDumb (NeurIPS 23’) 42.3 40.0 34.1 30.5 42.4 31.9 29.5 35.7 35.9 43.6 30.4 41.9 36.9 38.1 40.5 36.9
DeYO (ICLR 24’) 57.2 53.4 38.8 34.7 47.3 37.3 34.1 40.8 40.5 50.6 33.3 45.8 41.5 94.5 45.7 46.4
UnMix-TNS (ICLR 24’) 65.8 64.1 46.4 37.5 51.7 36.0 36.4 38.5 39.4 51.1 29.3 42.8 43.2 67.8 49.4 46.6
FreDA (ours) 34.8 34.7 36.6 29.4 41.2 29.9 28.4 33.8 33.7 41.1 29.8 34.9 36.9 37.1 38.7 34.7
IN-C (ResNet-50) 97.8 97.1 98.2 81.7 89.8 85.2 77.9 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
TBN 92.8 91.1 92.5 87.8 90.2 87.2 82.2 82.2 82.0 79.8 48.0 92.5 83.5 75.6 70.4 82.5
TENT (ICLR 21’) 99.2 98.7 99.0 90.5 95.1 90.5 84.6 86.6 84.0 86.5 46.7 98.1 86.1 77.7 72.9 86.4
ETA (ICML 22’) 90.7 89.2 90.5 77.0 80.6 74.0 68.9 72.4 70.3 64.6 43.9 93.4 69.2 52.3 55.9 72.9
AdaContrast (CVPR 22’) 96.2 95.5 96.2 93.2 96.4 96.3 90.5 92.7 91.9 92.4 50.8 97.0 96.6 89.7 87.1 90.8
CoTTA (CVPR 22’) 89.1 86.6 88.5 80.9 87.2 81.1 75.8 73.3 75.2 70.5 41.6 85.0 78.1 65.6 61.6 76.0
SAR (ICLR 23’) 98.4 97.3 98.0 84.0 87.3 82.6 77.2 77.5 76.1 72.5 43.1 96.0 78.3 61.8 60.4 79.4
RoTTA (CVPR 23’) 89.4 88.6 89.3 83.4 89.1 86.2 80.0 78.9 76.9 74.2 37.4 89.6 79.5 69.0 59.6 78.1
RDumb (NeurIPS 23’) 89.0 87.6 88.6 78.1 82.3 75.2 70.1 73.0 71.0 65.1 43.9 92.6 70.7 53.7 56.3 73.1
DeYO (ICLR 24’) 99.5 99.2 99.5 89.5 95.0 83.9 78.8 75.0 87.8 79.2 47.3 99.2 92.4 59.0 60.4 83.0
UnMix-TNS (ICLR 24’) 91.7 92.8 91.7 92.3 93.4 91.5 84.8 86.3 84.1 85.0 62.0 96.5 88.6 81.7 77.3 86.7
FreDA (ours) 72.4 74.0 71.4 76.5 82.3 72.1 64.1 64.4 64.8 59.1 43.7 79.7 71.0 54.2 58.6 67.2
VII Experiments
VII-A Datasets and Experimental Setup
VII-A1 Datasets
To provide a comprehensive evaluation of TTA deployment, we test models over multiple datasets under three different scenarios:
  • Common Image Corruptions: We evaluate models on CIFAR-10-C, CIFAR-100-C, and ImageNet-C [14] with 10, 100 and 1000 classes, respectively. These benchmarks are designed to assess the model robustness against various corruptions. Each dataset consists of 15 distinct corruptions across five severity levels, resulting in 150,000 at each severity for CIFAR-10-C/100-C, and 750,000 for ImageNet-C.
  • Natural Domain Shifts: We extend evaluation to DomainNet126 [35], which presents natural shifts across four domains (Real, Clipart, Painting, Sketch) encompassing 126 classes as a subset of the larger DomainNet dataset.
  • Medical Application: Models are further evaluated on Camelyon17 [36], comprising over 450,000 histopathological patches from lymph node sections for binary classification of normal and tumor tissue, with data originating from five distinct healthcare centers.
For corruption datasets, the model is pretrained on the clean dataset and the 15 corruptions are randomly mixed as the target distribution. We leverage the highest severity = 5 in all the experiments. In DomainNet126 and Camelyon17, one subdomain is selected as the source, and the others serve as mixed target distributions. All reported results are averaged over runs with fixed seeds (0, 1, and 2).
VII-A2 Baselines
We compare FreDA with 10 models, including TBN [31], TENT [3], CoTTA [5], ETA [6], SAR [7], AdaContrast [4], RoTTA [25], RDumb [9], DeYO [10], and UnMix-TNS [28]. TBN [31] re-estimates batch normalization statistics from test data. TENT [3] minimizes prediction entropy to optimize batch normalization. CoTTA [5] addresses long-term test-time adaptation in changing environments. ETA [6] and SAR [7] exclude unreliable and redundant samples during optimization. AdaContrast [4] utilizes contrastive learning to refine pseudo-labels and improve feature learning. RoTTA [25] presents a robust batch normalization scheme with a memory bank for category-balanced estimation. RDumb [9] leverages weighted entropy and periodically resets the model to its pretrained state to prevent collapse. DeYO [10] quantifies the impact of object-destructive transformations for sample selection and weighting. UnMix-TNS [28] introduces a test-time normalization layer for non-i.i.d. environments by decomposing BN statistics. For fair comparisons, we conduct experiments using the open source online TTA repository [37] [^1] https://github.com/mariodoebler/test-time-adaptation, which provides codes and configurations of state-of-the-art TTA methods.
VII-A3 Pretrained Models
We utilize models from RobustBench [38], including WildResNet-28 [39] for CIFAR-10-C and ResNeXt-29 [40] for CIFAR-100-C, both pretrained by [41]. For ImageNet-C, the pretrained ResNet-50 [42] is obtained from torchvision. For DomainNet126, pretrained ResNet-50 is sourced from AdaContrast [4], while for Camelyon17, we train a DenseNet-121 [43] from scratch to 100 epochs with other training specifications outlined in the Wilds benchmark [44].
VII-A4 Hyperparameter Configuration
The batch size is set to 200, 64, 128 and 32 for CIFAR-10/100-C, ImageNet-C, DomainNet126 and Camelyon17 following the previous methods. The SGD optimizer is used with learning rates adjusted to 0.01, 0.0001, 0.001 and 0.00005, respectively. The learning rate is proportionally decreased in the experiment studying the effect of batch size. The Kmeans Size is 512, Clutser Number is 4, Communication Interval is 10 across all the tasks. The perturbation magnitude $\alpha$ is fixed to 0.1 and the coefficient $\lambda$ in loss function is fixed to 0.5. Two threshold in Eq. 5 is set to the same value for corruption datasets and DomainNet126 following ETA [6]. While for Camelyon17, the class diversity related threshold is adjusted to 0.9 empirically.
TABLE III: Classification error rate ($\downarrow$) on DomainNet126 and Camelyon17 under Mixed Distribution Shifts.
DomainNet126
Methods Real Painting Clipart Sketch Avg.
Source 45.2 41.6 49.5 45.3 45.4
TBN 45.5 39.9 45.9 37.5 42.2
TENT (ICLR 21’) 42.2 37.8 44.7 37.5 40.6
ETA (ICML 22’) 41.1 37.3 43.4 36.4 39.5
SAR (ICLR 23’) 43.2 38.5 44.8 37.0 40.9
DeYO (ICLR 24’) 40.9 36.4 43.6 36.9 39.4
FreDA (ours) 40.2 36.1 40.0 33.6 37.5
Camelyon17
C1 C2 C3 C4 C5 Avg.
21.6 43.6 52.5 47.4 47.6 42.5
26.5 38.5 31.7 39.4 32.8 33.8
44.7 50.5 49.9 49.1 48.6 48.6
47.4 52.5 47.9 49.9 39.2 47.4
26.5 38.5 31.7 39.4 32.8 33.8
50.4 50.3 48.8 51.7 50.5 50.4
18.6 24.7 24.8 40.5 30.8 27.9
Figure 5: Averaged classification error rate (↓) on CIFAR-10-C/100-C and ImageNet-C with various batch size under Mixed Domains.
VII-B Main Results
VII-B1 FreDA improves across diverse distribution shifts
Our method consistently attains the lowest error rates across all evaluated datasets (see TABLE V and III). Notably, on the Camelyon17 dataset, FreDA reduced the error rate to 27.9%, outperforming the next best method by 5.9%. This significant improvement is particularly noTABLE where other approaches falter if compared with no training (TBN), meaning they struggle to adapt to the complex medical imaging data. By effectively handling high variability and intricate patterns in the data, FreDA maintains superior accuracy and adaptability.
VII-B2 FreDA remains stable under various batch size
To simulate deployment with constrained batch sizes, we evaluate models under both varying batch sizes and mixed distributions. In Figure 5, we present the results on CIFAR-10-C, CIFAR-100-C, and ImageNet-C using batch sizes ranging from 200 (64) down to 1. Unlike other methods that significantly degrade as batch size decreases – for example the error rate of DeYO increases from 27.7% to 89.8% when batch size drops from 200 to 1 on CIFAR-10-C – FreDA consistently maintains strong performance. This stability demonstrates FreDA’s robustness, making it highly suiTABLE for real-world applications where large batches is not always feasible.
TABLE IV: Ablation study of FreDA.
DT SS SA C10 C100 IN
44.1 82.5 86.4
24.8 54.2 81.2
29.6 37.5 71.0
39.4 71.7 92.9
24.3 36.3 69.4
27.7 36.2 65.9
24.4 50.2 77.7
22.9 34.7 67.2
VII-B3 FreDA enhances adaptation via synergistic designs
This section validates our designs by ablating three key modules – Decentralized Training (DT), Sample Selection (SS), and Sample Augmentation (SA). The baseline here leverages only the entropy loss. From TABLE VI, we have the following observations: 1) Implementing decentralized training and sample selection alone results in substantial improvements, reducing error rates dramatically across all datasets. 2) Sample augmentation alone has the possibility to increase error rates, suggesting that although this approach introduces useful variability, it may introduce unexpected noise under the absence of proper selection or decentralized training. 3) The combined approach delivers the best performance across all datasets, showing the synergistic effect of our different designs.
TABLE V: Classification error rate ($\downarrow$) on ImageNet-C under Mixed Distribution Shifts using ViT-Base backbone.
Baseline & Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
IN-C (VitBase-LN) 65.8 67.3 65.3 68.8 74.4 64.3 66.6 56.8 45.2 48.6 29.2 81.8 57.1 60.8 50.2 60.2
TENT (ICLR 21’) 60.6 60.4 59.6 63.6 67.8 57.1 61.2 55.0 48.8 47.4 28.6 66.7 53.9 50.4 44.4 55.0
ETA (ICML 22’) 59.3 57.8 57.9 58.8 62.8 52.5 58.2 51.0 46.4 44.2 28.8 58.3 51.1 46.9 41.9 51.7
AdaContrast (CVPR 22’) 64.8 63.4 63.3 72.8 76.6 73.7 74.6 67.7 48.0 89.6 30.2 93.2 60.8 57.3 46.3 65.5
CoTTA (CVPR 22’) 89.4 92.0 88.9 93.6 92.6 90.6 86.5 94.9 88.2 86.6 75.8 96.5 85.7 93.5 84.6 89.3
SAR (ICLR 23’) 58.9 57.6 57.6 59.4 63.6 53.0 58.5 52.3 47.1 45.4 28.3 61.6 51.4 47.4 42.0 52.3
RoTTA (CVPR 23’) 64.4 65.6 63.7 67.6 71.3 59.8 64.1 52.7 43.5 48.6 27.9 78.5 54.3 60.4 50.1 58.2
RDumb (NeurIPS 23’) 59.7 58.5 58.5 60.0 64.1 54.0 59.0 52.0 46.7 44.5 28.6 61.2 51.9 48.3 42.6 52.6
DeYO (ICLR 24’) 60.0 58.6 58.8 58.8 62.4 61.9 50.9 46.7 51.9 45.2 29.7 55.7 51.6 45.8 42.8 52.1
FreDA (ours) 55.9 53.7 55.0 58.0 57.9 50.9 57.4 45.5 42.9 43.9 29.5 51.7 47.8 41.6 40.7 48.8
VII-C More analysis and Discussion
VII-C1 Performance with Transformer Backbone
In addition to evaluating our model on commonly compared CNN backbones, we further assess its performance under transformer-based architecture, specifically using the ViT-Base backbone. Here, we report results on the ImageNet-C benchmark using ViTBase-LN [45] (see TABLE V), where the pretrained weights are obtained from torchvision. Importantly, all experimental configurations are kept consistent with those used in the CNN-based experiments. As shown, our method continues to deliver strong performance, demonstrating its robustness and adaptability across different backbone architectures.
TABLE VI: Sensitivity analysis on different datasets.
CLUSTER_NUM 2 4 8 16
CIFAR10-C 23.0 22.9 23.2 24.7
CIFAR100-C 34.8 34.7 34.7 35.6
IN-C (ResNet) 68.6 67.2 67.1 70.5
IN-C (ViT) 50.3 48.8 49.9 50.0
KMEANS_SIZE 256 512 1024 2048
CIFAR10-C 23.0 22.9 23.0 22.9
CIFAR100-C 34.6 34.7 34.8 34.8
IN-C (ResNet) 69.0 67.2 67.6 67.0
IN-C (ViT) 49.0 48.8 48.7 48.8
COMM_INTERVAL 1 10 100 1000
CIFAR10-C 22.6 22.9 22.6 22.0
CIFAR100-C 34.7 34.7 34.9 43.2
IN-C (ResNet) 67.1 67.2 67.2 67.4
IN-C (ViT) 48.4 48.8 48.8 48.7
PERT_MAGNITUDE 0.0 0.1 0.2 0.3
CIFAR10-C 24.3 22.9 22.5 22.2
CIFAR100-C 36.3 34.7 34.9 34.9
IN-C (ResNet) 69.3 67.2 67.0 66.9
IN-C (ViT) 49.6 48.8 48.5 48.6
VII-C2 Sensitivity Study
We investigate the impact of the key hyperparameters: CLUSTER_NUM, KMEANS_SIZE, COMM_INTERVAL, and PERT_MAGNITUDE. From TABLE VI, we have the following observations: 1) The choice of CLUSTER_NUM influences model performance, especially on more complex datasets. While performance remains relatively stable even with a small number of clusters (e.g., two clusters), increasing the number of clusters beyond four tends to lead to slight performance degradation, particularly on datasets like ImageNet-C. This suggests that while more clusters can improve adaptation capacity, excessively increasing the number can introduce overfitting and reduce generalization. Therefore, a moderate number of clusters, around 4, appears to strike a good balance between adaptation flexibility and maintaining model robustness. 2) Varying KMEANS_SIZE from 256 to 2048 results in stable performance across all datasets, indicating that our method is robust to changes in cluster sizes. 3) Our approach shows general robustness to communication frequency (varying COMM_INTERVAL from 1 to 1,000). On simpler datasets such as CIFAR10-C, infrequent communication (e.g., interval $f=1000$) performs best, likely due to effective independent learning. In contrast, for complex datasets, more frequent communication (e.g., $f=1$) improves performance, likely by mitigating divergence among local branches and ensuring model consistency. 4) Adjusting the perturbation magnitude (PERT_MAGNITUDE) has a noticeable effect, particularly on simpler datasets. Increasing perturbation improves performance, suggesting that spectral augmentation contributes to better generalization. However, for more complex datasets such as CIFAR100-C, higher perturbation levels lead to slight decline in performance, likely due to noise disrupting the alignment of finer features. A perturbation magnitude of 0.1 appears to strike an effective balance, offering the benefits of augmentation without negatively impacting feature alignment.
TABLE VII: Classification error rate ($\downarrow$) on CIFAR-10-C (C10), CIFAR-100-C (C100), and ImageNet-C (IN) using ResNet-50 & ViT-Base backbones under Continual Setting, averaged over 15 corruptions.
Methods C10 C100 IN(ResNet) IN(ViT)
Source 43.5 46.5 82.0 60.2
TBN 20.4 35.4 68.6 -
TENT (ICLR 21’) 20.0 62.2 62.6 54.5
ETA (ICML 22’) 17.9 32.2 60.2 49.8
AdaContrast (CVPR 22’) 18.5 33.5 65.5 57.0
CoTTA (CVPR 22’) 16.5 32.8 63.1 77.0
SAR (ICLR 23’) 20.4 32.0 61.9 51.7
RoTTA (CVPR 23’) 19.3 34.8 67.3 58.3
RDumb (NeurIPS 23’) 17.8 34.1 90.6 50.2
DeYO (ICLR 24’) 87.0 98.1 90.6 94.3
UnMix-TNS (ICLR 24’) 24.9 32.7 75.4 -
FreDA (ours) 19.5 32.5 60.2 47.9
VII-C3 Performance under continual settings
Although our method is specifically designed for mixed domain scenarios, we also evaluated its performance under the conventional continual test-time adaptation [5, 3] setting to assess its robustness in different contexts. In this setting, the model adapts online to a sequence of test domains without explicit knowledge of domain shifts, with only one distribution shift occurring at a time and not reappearing. Without adjusting any parameters, our method demonstrated competitive performance compared to current state-of-the-art approaches. Notably, while UnMix-TNS effectively addresses non-i.i.d. issues (dependent sampling at the class level), it is less effective under i.i.d. conditions. Our results suggest that the proposed FreDA not only excels in its intended mixed domain scenarios but also generalizes effectively to standard continual adaptation tasks, providing a robust solution across various distributional challenges.
VIII Conclusion and Future Work
This paper advances Test-Time Adaptation (TTA) by tackling the real-world complexities of heterogeneous data streams. Our decentralized approach precisely manages diverse data shifts, improving model adaptation in varied settings. By integrating Fourier-based augmentation, we expand the range of confident samples for each distribution shift, further boosting model performance. The experimental results underscore the efficacy of FreDA, highlighting its potential to influence the field and guide future research in adapting to dynamic and diverse data shifts. While FreDA addresses a critical challenge in handling heterogeneous data streams, providing a solid pipeline for this issue, there are still avenues for further enhancement. Our current aggregation approach, which averages models based on cluster counts, has been effective in solving the problem at hand. However, exploring alternative strategies – such as weighting models by the divergence between clusters – might lead to incremental improvements. Additionally, refining the sample selection process from a original sample-level focus to a more granular patch-level could extend FreDA’s applicability to tasks such as segmentation, further enhancing its versatility in real-world scenarios.
-A Adaptation Scenarios
-A1 Mixed Domains
: In this scenario, the model processes a long sequence of test samples where each sample $x_{i}\sim p_{e_{i}}(x)$ is independently drawn from a randomly selected target domain $\mathcal{D}_{e_{i}}\in\{\mathcal{D}_{t_{1}},\mathcal{D}_{t_{2}},\dots,\mathcal{D}_{t_{N}}\}$ and a randomly selected class $c_{i}$ among classes $\{1,2,\dots,C\}$. The sequence is represented as:
$\left\{x_{1}^{\mathcal{D}_{e_{1}},\,c_{1}},\ x_{2}^{\mathcal{D}_{e_{2}},\,c_{2}},\ \ \dots,\ x_{k}^{\mathcal{D}_{e_{k}},\,c_{k}}\right\},$
where each target domain index $e_{i}\in\{1,2,\dots,N\}$ and class number $c_{i}\in\{1,2,\dots,C\}$ are independently and randomly selected for each sample $x_{i}$.
-A2 Continual Domain Adaptation
: In this setting, the model adapts online to a sequentially presented series of test domains, where each domain shift occurs only once and does not reappear. The sequence progresses through distinct target domains in a fixed order $D_{1}\to D_{2}\to\dots\to D_{N}$, with samples from each domain appearing contiguously. The class labels within each domain are independently and randomly selected. The sequence is structured as:
$\begin{split}&\underbrace{\{(x_{1}^{D_{1},c_{1}}),(x_{2}^{D_{1},c_{2}}),\dots,(x_{l_{1}}^{D_{1},c_{l_{1}}})\}}_{\text{Samples from domain }D_{1}}\\ &\to\underbrace{\{(x_{l_{1}+1}^{D_{2},c_{l_{1}+1}}),\dots,(x_{l_{1}+l_{2}}^{D_{2},c_{l_{1}+l_{2}}})\}}_{\text{Samples from domain }D_{2}}\\ &\to\dots\to\underbrace{\{(x_{l_{1}+\dots+l_{N-1}+1}^{D_{N},c_{l_{1}+\dots+l_{N-1}+1}}),\dots,(x_{k}^{D_{N},c_{k}})\}}_{\text{Samples from domain }D_{N}}\end{split}$
where:
  • $D_{i}$ denotes the $i$-th target domain in the fixed sequence, with $i\in\{1,2,\dots,N\}$.
  • $c_{m}\in\{1,2,\dots,C\}$ is the randomly selected class label for the $m$-th sample, independent of domain transitions.
Figure 6: Examples from ImageNet-C under common image corruptions. The images showcase a range of corruption types (e.g., noise, blur, and weather distortions) at varying severity levels.
Figure 7: Samples from DomainNet126 across four subdomains (Real, Sketch, Painting, Clipart). These visualizations reflect the stylistic and perceptual variations inherent in each domain.
-B Dataset Visualization
To further illustrate the characteristics of the datasets used in our evaluation, we present visualizations of the data distribution across different corruption types (Fig. 6), natural domain shifts (Fig. 7), and medical centers (Fig. 2). These figures highlight the diverse challenges that our models face in each evaluation scenario, providing insight into the complexity of the test conditions.
-C Relation to Frequency Domain Learning
Recent advances highlight frequency-based techniques as powerful tools for domain transfer. In domain generalization, frequency analysis has revealed critical insights into model robustness and learning dynamics [46, 47, 48, 49, 50, 51, 52, 53]. For domain adaptation, interpolating image amplitude spectra across styles has proven effective in reducing domain gaps and preventing overfitting to low-level statistics [54, 55, 56, 57]. Motivated by these advancements, we discover that frequency information inherently captures domain characteristics, making it a valuable medium for decoupling mixed target domains – an aspect largely unexplored in prior work. On this basis, we propose FreDA, the first framework addressing TTA under mixed domain shifts in Fourier space, with a decentralized adaptation and perturbation mechanism.
-D Relation to Multi-Target Domain Adaptation
TTA under mixed distribution shifts shares similarities with multi-target unsupervised domain adaptation (MT-UDA) [11, 12, 58, 59], yet diverges critically in complexity: while MT-UDA assumes static, predefined target domains and leverages labeled source data for explicit domain alignment, TTA operates with no access to source data and must adapt to dynamic, unpredicTABLE target streams in an online manner. This eliminates direct source-target discrepancy computation and demand robust incremental adaptation rather than offline multi-domain optimization. These differences highlight the need for novel methodologies tailored specifically to TTA, going beyond the solutions developed for MT-UDA.
-E Relation to Decentralized, Federated, Distributed Learning
This work also intersects with decentralized, federated, and distributed learning due to splitting data batches into disjoint subsets and applying decentralized model adaptation. First, while decentralized learning focuses on non-i.i.d. data that is naturally distributed across multiple nodes [60], our approach starts with a centralized batch of target samples. By proactively splitting this data into disjoint subsets, we expose latent non-i.i.d. characteristics, enabling the effective use of decentralized learning techniques. Second, federated learning considers data privacy and model collaborations within decentralized learning [61]. In our case, as target samples are mixed in a batch, data privacy is not a concern. However, similar to federated learning, our approach also involves weight aggregation over subnetworks to enhance their base models before next batch adaptation. Third, distributed learning aims to improve training efficiency on large-scale datasets by partitioning data for synchronized training [62]. In contrast, our method operates in a real-time fine-tuning context with limited data at one time, hence scalability is less of a concern.

### v4

@@
 Deploying deep models in real-world scenarios remains challenging due to significant performance drops under distribution shifts between training and deployment environments.
-Test-Time Adaptation (TTA) has recently emerged as a promising solution, enabling on-the-fly model adaptation without access to source data.
-However, its effectiveness degrades significantly in the presence of complex, mixed distribution shifts – common in practical settings – where multiple latent domains coexist.
-Adapting under such intrinsic heterogeneity, especially in unlabeled and online conditions, remains an open and underexplored challenge.
-In this paper, we study TTA under mixed distribution shifts and move beyond conventional homogeneous adaptation paradigms.
-By revisiting TTA from a frequency-domain perspective, we observe that distribution heterogeneity often manifests in Fourier space – for instance, high-frequency components tend to carry domain-specific variations.
-This motivates us to perform domain-aware separation using high-frequency texture cues, making diverse shift patterns more tractable.
-To this end, we propose FreDA, a novel Frequency-based Decentralized Adaptation framework that decomposes globally heterogeneous data into locally homogeneous components in the frequency domain.
-It further employs decentralized learning and augmentation strategies to robustly adapt under complex, evolving shifts.
-Extensive experiments across various environments (corrupted, natural, and medical) demonstrate the superiority of our proposed framework over the state-of-the-arts.
-Index Terms:
-Test-time Adaptation, Transfer Learning
+Test-Time Adaptation (TTA) has recently emerged as a promising solution, enabling on-the-fly model adaptation.
+However, its effectiveness deteriorates in the presence of mixed distribution shifts – common in practical settings – where multiple target domains coexist.
+In this paper, we study TTA under mixed distribution shifts and move beyond conventional whole-batch adaptation paradigms.
+By revisiting distribution shifts from a spectral perspective, we find that the heterogeneity across latent domains is often pronounced in Fourier space.
+In particular, high-frequency components encode domain-specific variations, which facilitates clearer separation of samples from different distributions.
+Motivated by this observation, we propose to un-mix heterogeneous data streams using high-frequency domain cues, making diverse shift patterns more tractable.
+To this end, we propose Frequency-based Decentralized Adaptation (FreDA), a novel framework that decomposes globally heterogeneous data stream into locally homogeneous clusters in the Fourier space.
+It leverages decentralized learning and augmentation strategies to robustly adapt under mixed domain shifts.
+Extensive experiments across various environments (corrupted, natural, and medical) show the superiority of our method over the state-of-the-arts.
 I Introduction
@@
 It leverages the incoming data stream for real-time adjustments without the need for retraining on a labeled dataset, enabling swift model adaptation to unpredictable data characteristics during deployment.
-Despite their success, current TTA models often are limited to ideal testing conditions, typically involving homogeneous test samples with similar types of distribution shifts.
-In reality, distribution shifts are mixed, overlapping, and even conflicting [11, 12, 13, 14, 15].
+Despite their success, current TTA models are often limited to ideal testing conditions, typically involving homogeneous test samples with similar types of distribution shifts.
+In reality, distribution shifts are mixed, overlapping, and even conflicting [11, 12, 13, 14].
 For instance, photo management software handles diverse corruptions, such as noise, blur, compression; medical imaging platforms manage inconsistencies from varied acquisition methods; and autonomous driving systems face fluctuating conditions like lighting, weather, and road types.
 Such complexity poses serious challenges for existing methods.
-While recent efforts have extended TTA to continually changing environments [5], they typically assume a uniform target domain at each time step, as illustrated in the continual setting in Figure 1 (a).
+While recent efforts have extended TTA to continually changing environments [5], they typically assume a uniform target domain at each time step, as illustrated in the continual setting in Fig.
+1 (a).
 Some methods periodically reset the model to its source pre-trained state [7, 9] to mitigate the cumulative effect of sequential distribution shifts on adaptation.
 Others down-weight outlier target samples [6, 10] to suppress the impact of abrupt or rare shifts.
-Although these strategies aim to improve robustness under dynamic conditions, they fall short when facing the entangled / mixed real-world distributions (see Figure 1 (b)).
-Their shortcomings stem from a shared limitation: they lack the capacity to disentangle or localize domain variations.
-As a result, they often overfit to dominant patterns or suffer from catastrophic forgetting, failing to be deployed in realistic, heterogeneous data streams.
-Figure 1: (a) Illustration of three TTA scenarios with single, continual, and mixed domain shifts.
-(b) Classification error rate on CIFAR-10-C across different TTA settings: performance of conventional models drops sharply under mixed domain shifts.
-Capturing such intrinsic heterogeneity under unlabeled, online conditions is particularly challenging, and effective adaptation in this setting remains an open problem.
-To tackle this, we propose shifting from coarse, homogeneous adaptation to a fine-grained strategy that explicitly disentangles heterogeneous shifts.
-Our approach leverages the frequency information to disentangle domain variations, which naturally separates data features across frequency bands: high-frequency components capture fine details like edges and textures, while low-frequency components represent global structures such as shape and illumination (Section IV).
-By decomposing inputs accordingly, we better characterize distributional variations and identify diverse shifts.
-Moreover, the Fourier transform operates directly on raw pixel-level inputs, which shields the method from performance degradation under large domain gaps, and also enhancing adaptability and deployment efficiency in real-world applications.
-Building upon this insight, we introduce a framework, termed Frequency-based Decentralized Adaptation (FreDA).
-FreDA begins by partitioning incoming data in the Fourier domain, where high-frequency components are used to group samples with similar shift characteristics.
+Although these strategies aim to improve robustness under dynamic conditions, they lack the capacity to disentangle or localize domain variations within a batch.
+As a result, they often misalign conflicting domain shifts, failing to be deployed in realistic, heterogeneous data streams (see Fig.
+1 (b)).
+Figure 1: (a) Illustration of different online distribution shifts, ranging from single-domain to continual and mixed-domain shifts, with increasing complexity and realism.
+(b) Classification error rate on CIFAR-10-C: while conventional models experience a sharp performance drop as distribution shifts become more complex, our method maintains superior performance, particularly under mixed-domain shifts.
+To address this, we propose shifting from coarse, homogeneous / whole-batch adaptation to a fine-grained strategy that explicitly disentangles heterogeneous shifts.
+Our approach leverages the spectral information to disentangle domain variations, which naturally separates sample features across frequency bands: high-frequency components capture fine details like edges and textures, while low-frequency components represent global structures such as shape and illumination (see Section IV).
+By decomposing inputs accordingly, we are able to better characterize distributional variations and identify diverse shifts.
+Moreover, the Fourier transform operates directly on raw pixel-level inputs, independent of any pre-trained model, making it robust to large domain gaps and enhancing adaptability in real-world deployments.
+Figure 2: Illustration of FreDA framework.
+Heterogeneous data streams are partitioned in Fourier space into locally homogeneous clusters, where local models adapt independently and synchronize weights for intermittent update.
+Building upon this insight, we introduce Frequency-based Decentralized Adaptation (FreDA).
+FreDA begins by partitioning incoming data in the Fourier space, leveraging high-frequency components to group samples with similar domain characteristics.
 This transforms globally heterogeneous inputs into locally homogeneous subsets before any adaptation takes place.
-Based on these partitions, FreDA deploys multiple local models, each assigned to a specific subset.
-These models adapt independently to their respective data streams while periodically synchronizing through parameter exchange.
-The aggregated parameters form a shared base model, which is then used to reinitialize the local models, ensuring continual and coordinated adaptation across domains.
-This approach not only alleviates clustering errors via collaborative knowledge sharing but also allows local models to capture diverse distribution shifts.
-To further enhance robustness, we introduce a novel Fourier-based augmentation scheme that improves sample quality and strengthens adaptation to shift-specific characteristics.
+Multiple local models are then deployed, each assigned to a specific cluster, and adapt independently to their respective data streams while periodically uploading their parameters.
+These intermittent updates help build more stable model weights, reducing potential errors during clustering.
+To further enhance robustness, we introduce a Fourier-based augmentation scheme that improves sample quality and reinforces adaptation to shift-specific patterns.
 To summarize, the main contributions of this work are three-fold:
-  • We identify a key limitation of most existing TTA methods – their neglect of real-world data heterogeneity – which leads to suboptimal performance when confronted with mixed and diverse distribution shifts.
+  • We identify a key limitation of most existing TTA methods – their neglect of real-world data heterogeneity – which leads to suboptimal performance when confronted with diverse, mixed distribution shifts.
   • We propose FreDA, a frequency-based decentralized adaptation framework that leverages spectral decomposition and localized adaptation to effectively address heterogeneous distribution shifts at test-time.
-  • We validate FreDA through extensive experiments on corrupted, natural, and medical benchmarks, demonstrating consistent improvements over state-of-the-art methods across various TTA scenarios.
+  • We validate FreDA through extensive experiments on corrupted, natural, and medical benchmarks, demonstrating consistent improvements over state-of-the-art methods across various distribution shift scenarios.
 II Connections to Previous Studies
@@
 Deep neural networks often experience performance degradation when deployed in environments different from their training settings, due to distribution shifts between source (training) and target (testing) domains.
-This issue is extensively studied under the umbrella of transfer learning [16, 17], which aims to transfer knowledge across domains.
-A key branch of transfer learning is domain adaptation (DA) [18, 19, 20, 21], where models trained on a labeled source domain are adapted to a target domain.
-Depending on the availability of source and target data, DA can be classified into: Supervised DA using labeled data in both source and target domains during training; Unsupervised DA, which relies on labeled source data and unlabeled target data; and Source-free DA, that adapts to the target domain without access to source data, typically due to privacy or transmission constraints.
-TABLE I: Comparison of different transfer learning settings.
-Topic Source Data Target Labels Online Adaptation Model Agnostic
-Supervised Domain Adaptation $\checkmark$ $\checkmark$ $\times$
-Unsupervised Domain Adaptation $\checkmark$ $\times$ $\times$
-Source-free Domain Adaptation $\times$ $\times$ $\times$
-Test-time Training $\times$ $\times$ $\checkmark$ $\times$
-Test-time Adaptation $\times$ $\times$ $\checkmark$ $\checkmark$
-Note: TTA is the most practical setting for real-world deep model deployment.
-While traditional DA methods assume access to the whole target distribution during adaptation, such access is not always possible in real-world deployment.
-To address this, two prominent test-time adaptation paradigms have been proposed: Test-time Training (TTT) [22, 23, 24] introduces an auxiliary task – often self-supervised – during pre-training.
-At test time, the model optimizes this auxiliary objective to adapt to the target distribution accordingly; Test-time Adaptation (TTA) [3, 7, 8] poses a more demanding and practical scenario, where the model must adapt on-the-fly to the test stream without any prior modification during training.
+This issue has been extensively studied under the umbrella of transfer learning [15, 16, 17], which aims to transfer knowledge across domains.
+A key branch of transfer learning is domain adaptation (DA), where models trained on the labeled source domain are adapted to the target domain.
+Depending on the data availability, previous DA can be roughly classified into: Supervised DA [18] using labeled data in both source and target domains during training; Unsupervised DA [19, 20, 21, 22], which relies on labeled source data and unlabeled target data simultaneously; and Source-free DA [23, 24, 25], that adapts to the target domain without access to source data, typically due to privacy constraints.
+While traditional DA methods assume access to the entire target distribution before adaptation, real-world deployment often involves receiving data as an online stream, where samples arrive sequentially and the full distribution is never observed.
+This streaming nature of the target domain breaks the assumption of global access, making conventional domain adaptation infeasible.
+To handle this constraint, two prominent test-time adaptation paradigms have emerged.
+Online Test-time Training (OTTT) [26, 27] introduces an auxiliary task – often self-supervised – during pre-training.
+At test time, the model optimizes this auxiliary objective to adapt to the target distribution accordingly; Online Test-time Adaptation (OTTA) [3, 7, 6, 4, 5, 8] poses a more demanding and practical scenario, where the model must adapt on-the-fly to the test stream without any prior modification during training.
 This setting emphasizes the need for rapid, real-time model updates to effectively capture the continuously incoming data.
 A brief comparison of these adaptation settings is summarized in TABLE I.
-Our work falls under the TTA setting, where we propose a generalizable and robust adaptation framework that maintains strong performance across diverse distribution shifts and corruptions – without relying on source data or altering the pre-training process.
+Our work falls under the OTTA setting, where we propose a generalizable and robust adaptation framework that maintains strong performance across diverse distribution shifts and corruptions – without relying on source data or altering the pre-training process.
+TABLE I: Comparison of different transfer learning settings where $\checkmark$ denotes settings closer to real-world conditions.
+Research Fields No Source Data No Target Labels Online Adaptation Model Agnostic
+Supervised Domain Adaptation $\times$ $\times$ $\times$
+Unsupervised Domain Adaptation $\times$ $\checkmark$ $\times$
+Source-free Domain Adaptation $\checkmark$ $\checkmark$ $\times$
+Online Test-time Training $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+Online Test-time Adaptation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 II-B Non-i.i.d. Test-time Adaptation
@@
 II-B1 Mixed Domains
-Modern TTA methods designed for continual domain shifts [5, 25, 6, 9, 10] have pushed beyond the single adaptation setting.
-However, these approaches assume that the target domain is uniform at each time point – a special case of mixed domains where the change occurs across distinct time periods.
-In reality, data streams often involve genuinely mixed domains with heterogeneous distributions present simultaneously.
+Modern TTA methods designed for continual domain shifts [5, 28, 6, 9, 10] have pushed beyond single adaptation problems.
+However, while they account for the dynamic target distribution shifts over time, they typically assume that each batch contains only a single domain shift.
+This simplifies the real-world scenario, where multiple distribution shifts may occur simultaneously, creating a mixed-domain environment that remains largely unexplored.
 Existing approaches primarily focus on stabilizing model updates through periodic parameter resets [7, 9] or importance weighting of target samples [6, 9, 10], but these techniques are ill-equipped to disentangle and manage the complex interleaving of domain shifts, leading to degraded performance.
-In contrast, our work adopts a data-centric perspective by explicitly addressing mixed domains via frequency-space decomposition, which facilitates domain separation and proactive distribution alignment prior to model adaptation.
+In contrast, our work explicitly addresses mixed domains via frequency-space decomposition, which facilitates domain separation and proactive distribution alignment prior to model adaptation.
 While recent efforts like [7] acknowledge mixed distribution challenges under the broader “Dynamic Wild World” paradigm, their unified treatment of diverse real-world factors does not systematically address data heterogeneity.
 Conversely, our study directly targets the core issue of mixed domains in TTA, introducing tailored strategies for disentangling and adapting to co-occurring heterogeneous distributions.
-Figure 2: (a) Visual heterogeneity across five healthcare centers in Camelyon17 dataset: Example patches showcase diverse domain shifts that characterize mixed target domains.
-(b) Conventional sample latent features from pretrained models fail to separate different target subdomains, showing significant overlap.
-(c) High-frequency information enable distinct separation of target subdomains, contrasting with the conventional ones.
 II-B2 Dependent Sampling
-The second challenge stems from class-level dependencies introduced by temporal data sampling.
-This topic has garnered significant research attention, with a growing body of work [25, 26, 27, 28, 29, 30] actively addressing it through strategies such as pseudo-label-based rebalancing, or extended observation windows.
-These approaches aim to correct the temporal imbalance caused by skewed class distributions over time.
-However, unlike this these studies that focus on class-level imbalance, our work diverges by targeting sample-level heterogeneity, which remains underexplored.
-Specifically, we address style diversity and distributional variation at test-time – a challenge that persist even under class-balanced conditions – to improve model robustness under mixed-domain scenarios.
+The second challenge arises from temporal-dependent sampling, which introduces class-level dependencies.
+This topic has garnered considerable research attention, with a growing body of work [28, 29, 30, 31, 32, 33] actively addressing it through strategies such as pseudo-label-based rebalancing or extended observation windows.
+These approaches primarily aim to address temporal imbalances caused by skewed class distributions over time, aiming to improve target domain estimation by addressing low class diversity within a batch.
+However, as these methods focus on within-domain temporal coherence, their applicability becomes limited when batches contain samples from multiple target domains.
+This limitation is also reflected in our experiments, where mixed-domain sampling poses notable challenges not accounted for in these approaches.
 III Problem Definition of mixed domain shifts
@@
 TTA strategies must therefore refine the model to optimize its predictive accuracy across these diverse sub-domains, ensuring consistent and reliable performance amidst significant distributional variability.
-Methods C10 C100 IN
+IV Un-mixing TTA under Mixed Distribution Shifts: A Fourier Perspective
+While Test-Time Adaptation (TTA) methods perform well under single-type shifts, they degrade under mixed distribution shifts that are ubiquitous in real-world deployment (see Fig.
+1 (b)).
+To address this, we advocate a paradigm of proactive distribution management: before any parameter update, we resolve heterogeneity by decomposing the input stream in frequency space.
+Our key insight is that different distributions leave distinct spectral footprints, which can be utilized to disentangle the hetergenous data stream.
+As shown in Fig.
+3, high-frequency components naturally separate different domain shifts, in contrast to the substantial overlap observed in the latent sample features from a pre-trained model.
+This proactive step transforms an ill-posed, entangled adaptation problem into a set of well-conditioned homogeneous subproblems.
+Take TBN [34] as an example.
+In its original centralized version, TBN computes statistical information using the whole target batch and directly replaces the source values for online adaptation.
+While effective under single-domain shifts, this whole-batch, homogeneous strategy produces biased statistics under mixed domain shifts, leading to suboptimal adaptation (see Fig.
+4 (a)).
+In contrast, our decentralized approach first applies high-frequency clustering to partition the batch into locally homogeneous clusters.
+Batch normalization statistics are then computed within each cluster, enabling localized adaptation that can effectively prevent cross-domain contamination.
+Therefore, this decentralized strategy can reduce the error rate by a notable 4.27% and yields tighter within-class cohesion along with sharper between-class boundaries compared to centralized adaptation (see Fig.
+4 (b)).
+Figure 3: (a) Example histopathology patches from five distinct healthcare centers illustrate visual heterogeneity across different domains (Camelyon17 [35]).
+(b) Latent features from the pre-trained model fail to distinguish target subdomains and remain highly entangled.
+(c) In contrast, high-frequency features can clearly separate different target subdomains.
+V Frequency-based Decentralized Adaptation
+As outlined in the previous section, distinguishing samples from different distribution shifts provides a promising direction for TTA.
+Building on this insight, we introduce Frequency-based Decentralized Adaptation (FreDA), which partitions target samples in the Fourier space into homogeneous subdomains, enabling more accurate adaptation.
+We further enhance this process with a augmentation strategy that diversifies each subdomain in the Fourier space to improve model robustness.
+V-A Frequency-based Decentralized Learning
+We first introduce Frequency-based Decentralized Learning framework.
+It begins by extracting high-frequency information from the pixel space to partition target samples into homogeneous subsets, disclosing the latent target subdomains contained within a batch.
+We then perform a decentralized adaptation, where multiple local models – sharing the same architecture – independently adapt to their assigned subsets.
+At regular intervals, the parameters of these local models are aggregated to form a stronger base model, which is subsequently used to update the local models for the next adaptation phase.
+Detailed procedures are provided in the following subsections.
+Methods CIFAR-10-C CIFAR-100-C ImageNet-C
 TBN (centralized) 33.8 45.8 82.5
-TBN (Decentralized) 28.5 43.2 77.6
+TBN (our decentralized) 28.5 43.2 77.6
 (a) Centralized Adaptation
 (b) Decentralized Adaptation
-Figure 3: Experimental Result (error rate) and t-SNE feature visualization using TBN [31] as the pseudo-labeling method under mixed distribution shifts, comparing centralized and decentralized adaptation.
-Different colors represent different classes.
-(a) Centralized Adaptation: Global BN parameters are applied to the entire batch.
-(b) Decentralized Adaptation: Localized BN parameters are tailored for clusters separated based on high-frequency features.
-C10, C100, IN respectively denotes CIFAR-10-C, CIFAR-100-C and ImageNet-C datasets.
-IV TTA under Mixed Distribution Shifts: A Fourier Perspective
-While Test-Time Adaptation (TTA) methods excel under single-type distribution shifts, their performance degrades catastrophically when facing mixed distribution shifts – a ubiquitous challenge in real-world deployment.
-As demonstrated on CIFAR-10-C (Figure 1 (b)), leading methods like TENT [3], SAR [7] and RoTTA [25] suffer 16% average accuracy drop compared to single-shift scenarios.
-Conventional TTA paradigms attempt to mitigate this through passive robustness measures, such as entropy regularization [3], gradient masking [7], or model resetting [9].
-These methods treat mixed distribution shifts as irreducible noise, deferring corrective actions until model failures occur – a reactive paradigm that inevitably compromises adaptation efficacy.
-To address this, we propose a paradigm shift to proactive distribution management – instead of demanding models to handle arbitrary mixtures, we first resolve the heterogeneity via frequency-space decomposition.
-Our key insight is that mixed distributions exhibit spectral signatures that can be disentangled before model adaptation begins.
-As evidenced in Figure 2 (c), high-frequency components naturally separate distinct shifts without relying on error-prone model features (see Figure 2 (b).
-This theoretically grounded decomposition converts the ill-posed mixed-shift adaptation into multiple well-conditioned homogeneous-shift problems, effectively create an ideal adaptation environment from chaotic real-world streams.
-Crucially, this decomposition enables a decentralized adaptation where we can deploy domain-specific models tailored to high-frequency-separated clusters.
-Take TBN [31] as an example: while its centralized variant recalculates batch normalization (BN) statistics globally (treating heterogeneous data as a single distribution), our decentralized adaptation computes localized BN parameters within each spectrally disentangled cluster.
-As shown in Figu 3, this context-aware estimation preserves semantic discriminability: samples exhibit tighter within-class cohesion and sharper between-class boundaries compared to the blurred separability of centralized adaptation.
-Building on these insights, the following section details how leveraging the frequency domain enhances TTA methods for realistic scenarios with mixed distribution shifts.
-Figure 4: Illustration of Frequency-based Decentralized Learning.
-V Frequency-based Decentralized Adaptation
-The previous discussion highlights how heterogeneity within target distributions can hinder effective model adaptation.
-This naturally raises the question: How can we manage such distributional heterogeneity to enable better adaptation?
-As outlined earlier, distinguishing samples associated with different distribution shifts offers a promising pathway.
-Building on this insight, we address the TTA problem by leveraging high-frequency components in the data and propose a novel framework termed Frequency-based Decentralized Adaptation (FreDA).
-It partitions target samples into multiple homogeneous subdomains in the Fourier space, facilitating more accurate model adaptation.
-This is further enhanced by a frequency-based augmentation strategy that enriches each subdomain with diverse samples, strengthening model robustness.
-V-A Frequency-based Decentralized Learning
-Fourier transform offers an effective method to extract different frequency components from images, with high-frequency information particularly useful for capturing fine-grained details such as texture and noise.
-These details often highlight subtle variations among different distribution shifts.
-Based on this insight, we propose a module termed Frequency-based Decentralized Learning.
-It leverages frequency information extracted from the pixel space to systematically partition data into multiple homogeneous subsets, enabling multiple local models to specialize in capturing each distribution shift individually.
-Concurrently, our method enables collaborative learning by allowing periodic parameter sharing among these local models via an enhanced base model, thereby boosting overall adaptability to diverse distribution shifts.
+Figure 4: Classification error rate (%) and t-SNE sample feature visualization using TBN [34] under mixed distribution shifts.
+Different colors denote different classes.
+(a) Original Centralized Adaptation: global Batch Normalization parameters are computed from the whole target batch.
+(b) Our Decentralized Adaptation: local Batch Normalization parameters are tailored for clusters separated based on high-frequency features.
+Decentralized adaptation achieves a 4.27% error reduction and clearer class structure compared with centralized adaptation.
 V-A1 Frequency Feature Extraction
@@
 V-A4 Base Model Enhancement
-To integrate knowledge from all subnetworks and prevent degradation on specific subdomains, we periodically aggregation their parameters at intervals of time $T$:
+To mitigate potential clustering errors for more stable adaptation model weights, we allow each local models to periodically upload their parameters for aggregation at intervals of time $T$:
 $\mathbf{\theta}_{\text{base}}=\sum_{k=1}^{K}\left(\frac{|\mathcal{D}_{k}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}|}\mathbf{\theta}_{k}\right),$ (4)
 where $|\mathcal{D}_{k}|$ denotes sample number in cluster $k$.
-This aggregation step combines the parameter updates from each subnetwork proportionally to its cluster size.
-The updated parameters $\mathbf{\theta}_{\text{base}}$ are then distributed back to each subnetwork, initializing them for the next batch of training: $\mathbf{\theta}_{k}\leftarrow\mathbf{\theta}_{\text{base}}$.
+This aggregation step combines the parameters from all local models in proportion to their cluster size.
+The updated parameters $\mathbf{\theta}_{\text{base}}$ are then distributed back to each subnetwork, initializing them for the next adaptation round: $\mathbf{\theta}_{k}\leftarrow\mathbf{\theta}_{\text{base}}$.
 V-B Frequency-based Augmentation
-While decentralized learning effectively mitigates batch-level heterogeneity, it may falls short in accurately characterizing individual distribution shifts – largely due to limited batch size and noise introduced by coarse clustering.
-To improve target data quality, TTA methods commonly adopt data augmentation as a practical strategy to enhance model generalization.
-However, conventional augmentation techniques – typically borrowed from standard computer vision tasks such as rotation, cropping, and mixup – are primarily tailored for single-shift scenarios.
-They often fail to provide the targeted, distribution shifts-aware augmentation as required under mixed domain shifts.
-To this end, we propose a frequency-based augmentation strategy tailored for TTA in mixed-shift environments.
-Specifically, our method perturbs the amplitude components of target samples in the Fourier space, enabling targeted augmentation with respect to each sample’s underlying distribution.
-By manipulating the frequency-domain characteristics, our approach generates diverse yet distribution-consistent variations, thereby enriching subdomain representations and facilitating robust model adaptation
-Algorithm 1 Framework of Frequency-based Decentralized Learning and Augmentation
-1:Step $t$, Input batch $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}\in\mathbb{R}^{n\times h\times w\times c}$, Pretrained source model $q_{\theta}$, Initialize Feature Repository and Local Sample Pool $\mathcal{R},\mathcal{S}_{k}\leftarrow\emptyset$, CLUSTER_NUM $K$, KMEANS_SIZE $N$, COMM_INTERVAL $f$;
-2:Step 1: Extract Frequency Features
-3:for $i=1$ to $n$ do
-4:  $\mathbf{A}_{hf,i}\leftarrow\text{vec}(\text{G}(\mathbf{x}_{i}))$triangleright Extract high-freq components
-5:end for
-6:Step 2: Dynamic Clustering
-[…]
-[…]
-[…]
-10:Step 3: Local Model Training
-11:for cluster $k\in\{1,\ldots,K\}$ do
-12:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}\cup\{x_{i}\mid Z_{i}=k\}$triangleright Gather samples for cluster $k$
-13:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}[(|\mathcal{S}_{k}|-n+1):]$triangleright Keep the last $batch\_size=n$ entries
-14:  $\mathcal{S}_{k}^{\prime}\leftarrow\text{select\_samples}(\mathcal{S}_{k})$triangleright Select samples (Eq. 5)
-15:  for each $x_{i}\in\mathcal{S}_{k}^{\prime}$ do
-16:   $\tilde{x}_{i}\leftarrow\text{augment}(x_{i})$triangleright Augment data (Eq. 7)
-17:   $\text{Train}(q_{\theta_{k}},x_{i},\tilde{x}_{i})$triangleright Train local model (Eq. 3)
-18:  end for
-19:end for
-20:Step 4: Compile Predictions
-[…]
-22:Step 5: Base Model Enhancement
-23:If t % $f$ == 0 : $\triangleright$ Model Communication with interval $f$ (Eq.4)
-[…]
-[…]
-V-B1 Sample Selection Mechanism
+Although decentralized learning separates heterogeneous shifts into more coherent subdomains, each local model inevitably operates on a reduced set of samples.
+This limited within-cluster diversity may lead to unstable adaptation.
+To address this, we introduce a frequency-based augmentation strategy tailored for our decentralized paradigm.
+The method perturbs the amplitude components of target samples in the Fourier space, generating variations that are diverse yet consistent with each sample’s underlying distribution.
+This targeted augmentation enriches subdomain representations and strengthens model robustness across heterogeneous shifts.
+V-B1 Sample Selection Strategy
 We first select the reliable samples in each local model leveraging a criterion derived from the weighted entropy framework used in ETA [6] based on two primary conditions:
@@
 This ensures that selected samples exhibit significant deviations from previous predictions in class distribution and lower prediction uncertainty.
-V-B2 Frequency-Based Augmentation
+V-B2 Sample Augmentation Mechanism
 The augmentation process involves perturbing the amplitude spectrum.
@@
 $\tilde{\mathbf{X}}_{i}=\mathcal{F}^{-1}\left(\tilde{A}(x_{i}),P(x_{i})\right).$ (7)
+Algorithm 1 Framework of Frequency-based Decentralized Learning and Augmentation
+1:Step $t$, Input batch $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}\in\mathbb{R}^{n\times h\times w\times c}$, Pretrained source model $q_{\theta}$, Initialize Feature Repository and Local Sample Pool $\mathcal{R},\mathcal{S}_{k}\leftarrow\emptyset$, CLUSTER_NUM $K$, KMEANS_SIZE NN, COMM_INTERVAL $f$;
+2:Step 1: Extract Frequency Features
+3:for $i=1$ to $n$ do
+4:  $\mathbf{A}_{hf,i}\leftarrow\text{vec}(\text{G}(\mathbf{x}_{i}))$ $\triangleright$ Extract high-freq components
+5:end for
+6:Step 2: Dynamic Clustering
+[…]
+8:[$\mathcal{R}\leftarrow\mathcal{R}[(|\mathcal{R}|-N+1):]$ $\triangleright$ Keep the last NN entries for kmeans clustering
+[…]
+10:Step 3: Local Model Training
+11:for cluster $k\in\{1,\ldots,K\}$ do
+12:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}\cup\{x_{i}\mid Z_{i}=k\}$ $\triangleright$ Gather samples for cluster $k$
+13:  $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k}[(|\mathcal{S}_{k}|-n+1):]$ $\triangleright$ Keep the last batch_size=nbatch\_size=n entries
+14:  $\mathcal{S}_{k}^{\prime}\leftarrow\text{select\_samples}(\mathcal{S}_{k})$ $\triangleright$ Select samples (Eq. 5)
+15:  for each $x_{i}\in\mathcal{S}_{k}^{\prime}$ do
+16:   $\tilde{x}_{i}\leftarrow\text{augment}(x_{i})$ $\triangleright$ Augment data (Eq. 7)
+17:   $\text{Train}(q_{\theta_{k}},x_{i},\tilde{x}_{i})$ $\triangleright$ Train local model (Eq. 3)
+18:  end for
+19:end for
+20:Step 4: Compile Predictions
+[…]
+22:Step 5: Base Model Enhancement
+23:If t % $f$ == 0 : $\triangleright$ Model Communication with interval $f$ (Eq.4)
+24:  $\mathbf{\theta}_{\text{global}}\leftarrow\sum_{k=1}^{K}w_{k}\theta_{k}$
+25:  $\mathbf{\theta}_{k}\leftarrow\mathbf{\theta}_{\text{global}}$
 V-B3 Loss Function
@@
 During implementation, we leverage a memory bank strategy [28, 4, 36, 37] that is updated in real time.
-This design serves two purposes: 1) to ensure accurate clustering—since an overly small batch could impede the effective separation of data – and 2) to maintain the number of samples processed by the local model consistent with the original batch size, thereby preventing performance degradation due to a drastic reduction in batch size (e.g., reducing by a factor of the cluster number).
+This design serves two purposes: 1) to ensure accurate clustering – since an overly small batch could impede the effective separation of data – and 2) to maintain the number of samples processed by the local model consistent with the original batch size, thereby preventing performance degradation due to a drastic reduction in batch size (e.g., reducing by a factor of the cluster number).
 VI Theoretical Insights
-To facilitate theoretical understanding, we leverage the expansion-based analysis framework proposed in [34].
-We first revisit the core concepts and then highlight the theoretical advantages of our method under mixed domain shifts.
+To facilitate theoretical understanding of our method, we adopt the expansion-based framework [38], which offers a principled analysis of pseudo-labeling under local perturbations and neighborhood consistency, and connects these properties directly to upper bounds on model error rates.
+We first revisit the key definitions and theorems, then demonstrate how our method effectively tightens this generalization bound under mixed domain shifts from the viewpoints of Frequency-coherent Partition and Augmentation-induced Expansion.
 VI-A Expansion Theory
 Definition 1 ($(a,c)$-expansion).
-[…]
+A class-conditional distribution $P_{i}$ satisfies $(a,c)$-expansion if $\forall S\subseteq\mathcal{X}$ with $P_{i}(S)\leq a$:
 $\mathcal{P}_{i}(\mathcal{N}(S))\geq\min(c\mathcal{P}_{i}(S),1)$
@@
 $\mathrm{Err}(G)\leq\frac{2}{c-1}\mathrm{Err}(G_{pl})+\frac{2c}{c-1}\mu$
-where $\mathcal{L}_{pl}$ is pseudo-label loss and $\mathcal{R}_{B}$ is consistency regularizer.
+$\mathcal{L}_{pl}$ is pseudo-label loss and $\mathcal{R}_{B}$ is consistency regularizer.
 VI-B Analysis of Our Method
-Our method reduces both terms in Theorem 1 to achieve a tighter bound through two mechanisms:
-VI-B1 Frequency-Coherent Partitioning.
-Let $\{P_{t_{k}}\}_{k=1}^{K}$ be the $K$ sub-domains identified by frequency-based partitioning.
-For measurable sets $\{S_{k}\}_{k=1}^{K}$ where $S_{k}\subseteq X_{k}$ (the support of $P_{t_{k}}$), define:
+Our method reduces both terms in Theorem 1 to achieve a tighter bound from two perspectives:
+VI-B1 Frequency-based Partition
+Let $\{P_{t_{k}}\}_{k=1}^{K}$ be the $K$ sub-domains derived from frequency-based partitioning. For the related measurable sets $\{S_{k}\}_{k=1}^{K}$ ($S_{k}\subseteq X_{k}$), we have:
 $\mathrm{Err}(G_{pl}^{k})\leq\mathrm{Err}(G_{pl})-\Delta_{k}\quad$
 where $\Delta_{k}=\mathcal{P}_{t_{k}}\left(E_{c}\right)$ and $E_{c}=\{x\in X_{k}\mid G_{pl}(x)\neq y(x)\}$.
-This contributes to the reduction of the first term in Theorem 1, which is further supported by experimental results in Figure 3.
+This contributes to the reduction of the first term in Theorem 1, which is further supported by experimental results in Fig.
+4.
 VI-B2 Augmentation-Induced Expansion
@@
 The expansion gain becomes:
-$\begin{split}\widehat{c}&=c_{\text{base}}+\gamma,\\ \gamma=\inf_{S:P(S)\leq a}&\frac{P(\widehat{\mathcal{N}}(S)\setminus\mathcal{N}_{\text{base}}(S))}{P(S)}\end{split}$
+$\widehat{c}=c_{\text{base}}+\gamma,\ \gamma=\inf_{S:P(S)\leq a}\frac{P(\widehat{\mathcal{N}}(S)\setminus\mathcal{N}_{\text{base}}(S))}{P(S)}$
 Substituting into Theorem 1, the error becomes:
 $\begin{split}\mathrm{Err}(G)&\leq\underbrace{\frac{2}{(c_{\text{base}}+\gamma)-1}\left(\mathrm{Err}(G_{pl})-\sum_{k=1}^{K}\Delta_{k}\right)}_{\text{Reduced by both terms}}\\ &+\underbrace{\frac{2(c_{\text{base}}+\gamma)}{(c_{\text{base}}+\gamma)-1}\mu}_{\text{Reduced by }O(\gamma/c^{2})}\end{split}$
-Notation Summary
+Remark.
+By leveraging frequency-based decentralized learning and augmentation, our method addresses the key components in the expansion-theoretic error, yielding a provably tighter generalization bound that supports more challenging online adaptation scenarios under diverse distribution shifts.
+Notation
   • $c$: Expansion factor (Def. 1).
@@
     Neighborhood enhancement: Additional expansion capability measured as the relative increase of augmented neighborhoods.
-TABLE II: Classification error rate ($\downarrow$) on CIFAR-10-C, CIFAR-100-C, and ImageNet-C (IN-C) under Mixed Distribution Shifts.
-Baseline & Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
-CIFAR-10-C (WRN-28) 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.4 30.3 43.5
+TABLE II: Classification error rate ($\downarrow$) on CIFAR-10-C, CIFAR-100-C, ImageNet-C under Mixed Distribution Shifts.
+Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
+CIFAR-10-C 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.4 30.3 43.5
 TBN 45.5 42.8 59.7 34.2 44.3 29.8 32.0 19.8 21.1 21.5 9.3 27.9 33.1 55.5 30.8 33.8
-TENT (ICLR 21’) 73.5 70.1 81.4 31.6 60.3 29.6 28.5 30.8 35.3 25.7 13.6 44.2 32.6 70.2 34.9 44.1
-ETA (ICML 22’) 36.2 33.3 52.3 22.9 38.9 22.4 20.5 19.5 19.7 20.4 11.3 35.4 26.6 38.8 25.1 28.2
-AdaContrast (CVPR 22’) 36.7 34.3 48.8 18.2 39.1 21.1 17.7 18.6 18.3 16.8 9.0 17.4 27.7 44.8 24.9 26.2
-CoTTA (CVPR 22’) 38.7 36.0 56.1 36.0 36.8 32.3 31.0 19.9 17.6 27.2 11.7 52.6 30.5 35.8 25.7 32.5
-SAR (ICLR 23’) 45.5 42.7 59.6 34.1 44.3 29.7 31.9 19.8 21.1 21.5 9.3 27.8 33.0 55.4 30.8 33.8
-RoTTA (CVPR 23’) 60.0 55.5 70.0 23.8 44.1 20.7 21.3 20.2 22.7 16.0 9.4 22.7 27.0 58.6 29.2 33.4
-RDumb (NeurIPS 23’) 34.9 32.3 49.4 23.3 38.2 23.3 20.7 19.9 19.3 20.7 11.2 29.3 26.7 41.5 25.2 27.7
-DeYO (ICLR 24’) 45.8 42.3 65.7 21.3 41.8 25.1 19.5 21.1 19.6 19.2 12.3 21.8 28.5 39.3 28.0 30.1
-UnMix-TNS (ICLR 24’) 50.0 44.4 44.3 34.4 48.2 32.7 30.0 35.5 35.9 47.5 28.1 38.7 43.9 40.0 43.3 39.8
+TENT 73.5 70.1 81.4 31.6 60.3 29.6 28.5 30.8 35.3 25.7 13.6 44.2 32.6 70.2 34.9 44.1
+ETA 36.2 33.3 52.3 22.9 38.9 22.4 20.5 19.5 19.7 20.4 11.3 35.4 26.6 38.8 25.1 28.2
+AdaContrast 36.7 34.3 48.8 18.2 39.1 21.1 17.7 18.6 18.3 16.8 9.0 17.4 27.7 44.8 24.9 26.2
+CoTTA 38.7 36.0 56.1 36.0 36.8 32.3 31.0 19.9 17.6 27.2 11.7 52.6 30.5 35.8 25.7 32.5
+SAR 45.5 42.7 59.6 34.1 44.3 29.7 31.9 19.8 21.1 21.5 9.3 27.8 33.0 55.4 30.8 33.8
+RoTTA 60.0 55.5 70.0 23.8 44.1 20.7 21.3 20.2 22.7 16.0 9.4 22.7 27.0 58.6 29.2 33.4
+RDumb 34.9 32.3 49.4 23.3 38.2 23.3 20.7 19.9 19.3 20.7 11.2 29.3 26.7 41.5 25.2 27.7
+DeYO 45.8 42.3 65.7 21.3 41.8 25.1 19.5 21.1 19.6 19.2 12.3 21.8 28.5 39.3 28.0 30.1
+UnMix-TNS 50.0 44.4 44.3 34.4 48.2 32.7 30.0 35.5 35.9 47.5 28.1 38.7 43.9 40.0 43.3 39.8
 FreDA (ours) 23.1 22.2 32.2 18.7 41.6 18.8 16.8 17.9 19.9 16.9 9.8 13.2 29.1 35.4 28.6 22.9
-CIFAR-100-C (ResNeXt-29) 73.0 68.0 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
+CIFAR-100-C 73.0 68.0 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
 TBN 62.7 60.7 43.1 35.5 50.3 35.7 34.4 39.9 51.5 27.5 45.5 42.3 72.8 46.4 45.8 45.8
-TENT (ICLR 21’) 95.6 95.2 89.2 72.8 82.9 74.4 72.3 78.0 79.7 84.7 71.0 88.5 77.8 96.8 78.7 82.5
-ETA (ICML 22’) 42.6 40.3 34.1 30.3 42.4 32.0 29.4 35.6 35.8 44.1 30.2 41.8 36.9 38.9 40.9 37.0
-AdaContrast (CVPR 22’) 54.5 51.5 37.6 30.7 45.4 32.1 30.3 36.9 36.5 45.3 28.0 42.7 38.2 75.4 41.7 41.8
-CoTTA (CVPR 22’) 54.4 52.7 49.8 36.0 45.8 36.7 33.9 38.9 35.8 52.0 30.4 60.9 40.2 38.0 41.1 43.1
-SAR (ICLR 23’) 75.8 72.7 41.1 29.2 45.2 31.1 28.9 36.7 37.7 43.9 29.3 41.8 37.1 89.2 42.4 45.5
-RoTTA (CVPR 23’) 65.0 62.3 39.3 33.4 50.0 34.2 32.6 36.6 36.5 45.0 26.4 41.6 40.6 89.5 48.5 45.4
-RDumb (NeurIPS 23’) 42.3 40.0 34.1 30.5 42.4 31.9 29.5 35.7 35.9 43.6 30.4 41.9 36.9 38.1 40.5 36.9
-DeYO (ICLR 24’) 57.2 53.4 38.8 34.7 47.3 37.3 34.1 40.8 40.5 50.6 33.3 45.8 41.5 94.5 45.7 46.4
-UnMix-TNS (ICLR 24’) 65.8 64.1 46.4 37.5 51.7 36.0 36.4 38.5 39.4 51.1 29.3 42.8 43.2 67.8 49.4 46.6
+TENT 95.6 95.2 89.2 72.8 82.9 74.4 72.3 78.0 79.7 84.7 71.0 88.5 77.8 96.8 78.7 82.5
+ETA 42.6 40.3 34.1 30.3 42.4 32.0 29.4 35.6 35.8 44.1 30.2 41.8 36.9 38.9 40.9 37.0
+AdaContrast 54.5 51.5 37.6 30.7 45.4 32.1 30.3 36.9 36.5 45.3 28.0 42.7 38.2 75.4 41.7 41.8
+CoTTA 54.4 52.7 49.8 36.0 45.8 36.7 33.9 38.9 35.8 52.0 30.4 60.9 40.2 38.0 41.1 43.1
+SAR 75.8 72.7 41.1 29.2 45.2 31.1 28.9 36.7 37.7 43.9 29.3 41.8 37.1 89.2 42.4 45.5
+RoTTA 65.0 62.3 39.3 33.4 50.0 34.2 32.6 36.6 36.5 45.0 26.4 41.6 40.6 89.5 48.5 45.4
+RDumb 42.3 40.0 34.1 30.5 42.4 31.9 29.5 35.7 35.9 43.6 30.4 41.9 36.9 38.1 40.5 36.9
+DeYO 57.2 53.4 38.8 34.7 47.3 37.3 34.1 40.8 40.5 50.6 33.3 45.8 41.5 94.5 45.7 46.4
+UnMix-TNS 65.8 64.1 46.4 37.5 51.7 36.0 36.4 38.5 39.4 51.1 29.3 42.8 43.2 67.8 49.4 46.6
 FreDA (ours) 34.8 34.7 36.6 29.4 41.2 29.9 28.4 33.8 33.7 41.1 29.8 34.9 36.9 37.1 38.7 34.7
-IN-C (ResNet-50) 97.8 97.1 98.2 81.7 89.8 85.2 77.9 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
+ImageNet-C 97.8 97.1 98.2 81.7 89.8 85.2 77.9 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
 TBN 92.8 91.1 92.5 87.8 90.2 87.2 82.2 82.2 82.0 79.8 48.0 92.5 83.5 75.6 70.4 82.5
-TENT (ICLR 21’) 99.2 98.7 99.0 90.5 95.1 90.5 84.6 86.6 84.0 86.5 46.7 98.1 86.1 77.7 72.9 86.4
-ETA (ICML 22’) 90.7 89.2 90.5 77.0 80.6 74.0 68.9 72.4 70.3 64.6 43.9 93.4 69.2 52.3 55.9 72.9
-AdaContrast (CVPR 22’) 96.2 95.5 96.2 93.2 96.4 96.3 90.5 92.7 91.9 92.4 50.8 97.0 96.6 89.7 87.1 90.8
-CoTTA (CVPR 22’) 89.1 86.6 88.5 80.9 87.2 81.1 75.8 73.3 75.2 70.5 41.6 85.0 78.1 65.6 61.6 76.0
-SAR (ICLR 23’) 98.4 97.3 98.0 84.0 87.3 82.6 77.2 77.5 76.1 72.5 43.1 96.0 78.3 61.8 60.4 79.4
-RoTTA (CVPR 23’) 89.4 88.6 89.3 83.4 89.1 86.2 80.0 78.9 76.9 74.2 37.4 89.6 79.5 69.0 59.6 78.1
-RDumb (NeurIPS 23’) 89.0 87.6 88.6 78.1 82.3 75.2 70.1 73.0 71.0 65.1 43.9 92.6 70.7 53.7 56.3 73.1
-DeYO (ICLR 24’) 99.5 99.2 99.5 89.5 95.0 83.9 78.8 75.0 87.8 79.2 47.3 99.2 92.4 59.0 60.4 83.0
-UnMix-TNS (ICLR 24’) 91.7 92.8 91.7 92.3 93.4 91.5 84.8 86.3 84.1 85.0 62.0 96.5 88.6 81.7 77.3 86.7
+TENT 99.2 98.7 99.0 90.5 95.1 90.5 84.6 86.6 84.0 86.5 46.7 98.1 86.1 77.7 72.9 86.4
+ETA 90.7 89.2 90.5 77.0 80.6 74.0 68.9 72.4 70.3 64.6 43.9 93.4 69.2 52.3 55.9 72.9
+AdaContrast 96.2 95.5 96.2 93.2 96.4 96.3 90.5 92.7 91.9 92.4 50.8 97.0 96.6 89.7 87.1 90.8
+CoTTA 89.1 86.6 88.5 80.9 87.2 81.1 75.8 73.3 75.2 70.5 41.6 85.0 78.1 65.6 61.6 76.0
+SAR 98.4 97.3 98.0 84.0 87.3 82.6 77.2 77.5 76.1 72.5 43.1 96.0 78.3 61.8 60.4 79.4
+RoTTA 89.4 88.6 89.3 83.4 89.1 86.2 80.0 78.9 76.9 74.2 37.4 89.6 79.5 69.0 59.6 78.1
+RDumb 89.0 87.6 88.6 78.1 82.3 75.2 70.1 73.0 71.0 65.1 43.9 92.6 70.7 53.7 56.3 73.1
+DeYO 99.5 99.2 99.5 89.5 95.0 83.9 78.8 75.0 87.8 79.2 47.3 99.2 92.4 59.0 60.4 83.0
+UnMix-TNS 91.7 92.8 91.7 92.3 93.4 91.5 84.8 86.3 84.1 85.0 62.0 96.5 88.6 81.7 77.3 86.7
 FreDA (ours) 72.4 74.0 71.4 76.5 82.3 72.1 64.1 64.4 64.8 59.1 43.7 79.7 71.0 54.2 58.6 67.2
+Note: WRN-28, ResNeXt-29, and ResNet-50 backbones are used for CIFAR-10-C, CIFAR-100-C, and ImageNet-C, respectively.
 VII Experiments
@@
 VII-A1 Datasets
-To provide a comprehensive evaluation of TTA deployment, we test models over multiple datasets under three different scenarios:
+To provide a comprehensive evaluation of TTA deployment, we test models under three different scenarios (see Fig. 3 and Appendix VIII-C for visualizations):
   • Common Image Corruptions: We evaluate models on CIFAR-10-C, CIFAR-100-C, and ImageNet-C [39] with 10, 100 and 1000 classes, respectively.
@@
 For fair comparisons, we conduct experiments using the open source online TTA repository [41] [^1] https://github.com/mariodoebler/test-time-adaptation, which provides codes and configurations of state-of-the-art TTA methods.
-VII-A3 Pretrained Models
-We utilize models from RobustBench [38], including WildResNet-28 [39] for CIFAR-10-C and ResNeXt-29 [40] for CIFAR-100-C, both pretrained by [41].
-For ImageNet-C, the pretrained ResNet-50 [42] is obtained from torchvision.
-For DomainNet126, pretrained ResNet-50 is sourced from AdaContrast [4], while for Camelyon17, we train a DenseNet-121 [43] from scratch to 100 epochs with other training specifications outlined in the Wilds benchmark [44].
-VII-A4 Hyperparameter Configuration
-The batch size is set to 200, 64, 128 and 32 for CIFAR-10/100-C, ImageNet-C, DomainNet126 and Camelyon17 following the previous methods.
+VII-A3 Hyperparameter Configuration
+The adaptation batch size is set to 200, 64, 128 and 32 for CIFAR-10/100-C, ImageNet-C, DomainNet126 and Camelyon17 following the previous methods.
 The SGD optimizer is used with learning rates adjusted to 0.01, 0.0001, 0.001 and 0.00005, respectively.
@@
 TBN 45.5 39.9 45.9 37.5 42.2
-TENT (ICLR 21’) 42.2 37.8 44.7 37.5 40.6
-ETA (ICML 22’) 41.1 37.3 43.4 36.4 39.5
-SAR (ICLR 23’) 43.2 38.5 44.8 37.0 40.9
-DeYO (ICLR 24’) 40.9 36.4 43.6 36.9 39.4
+TENT 42.2 37.8 44.7 37.5 40.6
+ETA 41.1 37.3 43.4 36.4 39.5
+SAR 43.2 38.5 44.8 37.0 40.9
+DeYO 40.9 36.4 43.6 36.9 39.4
 FreDA (ours) 40.2 36.1 40.0 33.6 37.5
@@
 18.6 24.7 24.8 40.5 30.8 27.9
-Figure 5: Averaged classification error rate (↓) on CIFAR-10-C/100-C and ImageNet-C with various batch size under Mixed Domains.
+Figure 5: Classification error rate (↓) on CIFAR-10-C/100-C with various batch size under Mixed Distribution Shifts.
+VII-A4 Implementation Details
+We utilize models from RobustBench [42], including WildResNet-28 [43] for CIFAR-10-C and ResNeXt-29 [44] for CIFAR-100-C, both pretrained by [45].
+For ImageNet-C, the pretrained ResNet-50 [46] is obtained from torchvision.
+For DomainNet126, pretrained ResNet-50 is sourced from AdaContrast [4], while for Camelyon17, we train a DenseNet-121 [47] from scratch to 100 epochs with other training specifications outlined in the Wilds benchmark [48].
+Our implementation can be found at https://github.com/kiwi12138/FreDA.
 VII-B Main Results
 VII-B1 FreDA improves across diverse distribution shifts
-Our method consistently attains the lowest error rates across all evaluated datasets (see TABLE V and III).
+Our method consistently attains the lowest error rates across all evaluated datasets (see TABLE II and III).
 Notably, on the Camelyon17 dataset, FreDA reduced the error rate to 27.9%, outperforming the next best method by 5.9%.
@@
 To simulate deployment with constrained batch sizes, we evaluate models under both varying batch sizes and mixed distributions.
-In Figure 5, we present the results on CIFAR-10-C, CIFAR-100-C, and ImageNet-C using batch sizes ranging from 200 (64) down to 1.
+In Fig.
+5, we present the results on CIFAR-10-C and CIFAR-100-C using batch sizes ranging from 200 down to 1.
 Unlike other methods that significantly degrade as batch size decreases – for example the error rate of DeYO increases from 27.7% to 89.8% when batch size drops from 200 to 1 on CIFAR-10-C – FreDA consistently maintains strong performance.
 This stability demonstrates FreDA’s robustness, making it highly suitable for real-world applications where large batches is not always feasible.
-TABLE IV: Ablation study of FreDA.
-DT SS SA C10 C100 IN
-44.1 82.5 86.4
-24.8 54.2 81.2
-29.6 37.5 71.0
-39.4 71.7 92.9
-24.3 36.3 69.4
-27.7 36.2 65.9
-24.4 50.2 77.7
-22.9 34.7 67.2
-VII-B3 FreDA enhances adaptation via synergistic designs
-This section validates our designs by ablating three key modules – Decentralized Training (DT), Sample Selection (SS), and Sample Augmentation (SA).
-The baseline here leverages only the entropy loss.
-From TABLE VI, we have the following observations: 1) Implementing decentralized training and sample selection alone results in substantial improvements, reducing error rates dramatically across all datasets.
-2) Sample augmentation alone has the possibility to increase error rates, suggesting that although this approach introduces useful variability, it may introduce unexpected noise under the absence of proper selection or decentralized training.
-3) The combined approach delivers the best performance across all datasets, showing the synergistic effect of our different designs.
-TABLE V: Classification error rate ($\downarrow$) on ImageNet-C under Mixed Distribution Shifts using ViT-Base backbone.
-Baseline & Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
+VII-B3 FreDA enhances adaptation via frequency-based designs
+This section validates our designs by ablating two key modules – Frequency-based Decentralized learning (FD) and Frequency-based Augmentation (FA) – from the full model.
+It is worth mentioning that our FreDA is built upon TENT (our baseline) by integrating both FD and FA.
+From Fig.
+6, we observe the following: a) Removing FD leads to a clear performance drop, underscoring the importance of our proposed decentralized adaptation paradigm in handling mixed distribution shifts.
+b) Removing FA results in a surprisingly large error rate increase on CIFAR-100-C.
+In more challenging classification scenarios (e.g., CIFAR-100-C vs.
+CIFAR-10-C), the target domain typically exhibits sparser class-wise coverage.
+Augmentation enriches the target sample diversity while preserving semantic consistency, thereby mitigating overfitting to spurious patterns and enhancing robustness under severe and complex distribution shifts.
+Nevertheless, even without FA, our method still significantly outperforms the baseline model, further confirming the critical role of our decentralized adaptation framework.
+Figure 6: Classification error rate (↓) on CIFAR-10-C/100-C with various batch size under Mixed Distribution Shifts.
+TABLE IV: Classification error rate ($\downarrow$) on ImageNet-C under Mixed Distribution Shifts using ViT-Base backbone.
+Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
 IN-C (VitBase-LN) 65.8 67.3 65.3 68.8 74.4 64.3 66.6 56.8 45.2 48.6 29.2 81.8 57.1 60.8 50.2 60.2
-TENT (ICLR 21’) 60.6 60.4 59.6 63.6 67.8 57.1 61.2 55.0 48.8 47.4 28.6 66.7 53.9 50.4 44.4 55.0
-ETA (ICML 22’) 59.3 57.8 57.9 58.8 62.8 52.5 58.2 51.0 46.4 44.2 28.8 58.3 51.1 46.9 41.9 51.7
-AdaContrast (CVPR 22’) 64.8 63.4 63.3 72.8 76.6 73.7 74.6 67.7 48.0 89.6 30.2 93.2 60.8 57.3 46.3 65.5
-CoTTA (CVPR 22’) 89.4 92.0 88.9 93.6 92.6 90.6 86.5 94.9 88.2 86.6 75.8 96.5 85.7 93.5 84.6 89.3
-SAR (ICLR 23’) 58.9 57.6 57.6 59.4 63.6 53.0 58.5 52.3 47.1 45.4 28.3 61.6 51.4 47.4 42.0 52.3
-RoTTA (CVPR 23’) 64.4 65.6 63.7 67.6 71.3 59.8 64.1 52.7 43.5 48.6 27.9 78.5 54.3 60.4 50.1 58.2
-RDumb (NeurIPS 23’) 59.7 58.5 58.5 60.0 64.1 54.0 59.0 52.0 46.7 44.5 28.6 61.2 51.9 48.3 42.6 52.6
-DeYO (ICLR 24’) 60.0 58.6 58.8 58.8 62.4 61.9 50.9 46.7 51.9 45.2 29.7 55.7 51.6 45.8 42.8 52.1
+TENT 60.6 60.4 59.6 63.6 67.8 57.1 61.2 55.0 48.8 47.4 28.6 66.7 53.9 50.4 44.4 55.0
+ETA 59.3 57.8 57.9 58.8 62.8 52.5 58.2 51.0 46.4 44.2 28.8 58.3 51.1 46.9 41.9 51.7
+AdaContrast 64.8 63.4 63.3 72.8 76.6 73.7 74.6 67.7 48.0 89.6 30.2 93.2 60.8 57.3 46.3 65.5
+CoTTA 89.4 92.0 88.9 93.6 92.6 90.6 86.5 94.9 88.2 86.6 75.8 96.5 85.7 93.5 84.6 89.3
+SAR 58.9 57.6 57.6 59.4 63.6 53.0 58.5 52.3 47.1 45.4 28.3 61.6 51.4 47.4 42.0 52.3
+RoTTA 64.4 65.6 63.7 67.6 71.3 59.8 64.1 52.7 43.5 48.6 27.9 78.5 54.3 60.4 50.1 58.2
+RDumb 59.7 58.5 58.5 60.0 64.1 54.0 59.0 52.0 46.7 44.5 28.6 61.2 51.9 48.3 42.6 52.6
+DeYO 60.0 58.6 58.8 58.8 62.4 61.9 50.9 46.7 51.9 45.2 29.7 55.7 51.6 45.8 42.8 52.1
 FreDA (ours) 55.9 53.7 55.0 58.0 57.9 50.9 57.4 45.5 42.9 43.9 29.5 51.7 47.8 41.6 40.7 48.8
-VII-C More analysis and Discussion
-VII-C1 Performance with Transformer Backbone
-In addition to evaluating our model on commonly compared CNN backbones, we further assess its performance under transformer-based architecture, specifically using the ViT-Base backbone.
-Here, we report results on the ImageNet-C benchmark using ViTBase-LN [45] (see TABLE V), where the pretrained weights are obtained from torchvision.
-Importantly, all experimental configurations are kept consistent with those used in the CNN-based experiments.
-As shown, our method continues to deliver strong performance, demonstrating its robustness and adaptability across different backbone architectures.
-TABLE VI: Sensitivity analysis on different datasets.
-CLUSTER_NUM 2 4 8 16
-CIFAR10-C 23.0 22.9 23.2 24.7
-CIFAR100-C 34.8 34.7 34.7 35.6
-IN-C (ResNet) 68.6 67.2 67.1 70.5
-IN-C (ViT) 50.3 48.8 49.9 50.0
-KMEANS_SIZE 256 512 1024 2048
-CIFAR10-C 23.0 22.9 23.0 22.9
-CIFAR100-C 34.6 34.7 34.8 34.8
-IN-C (ResNet) 69.0 67.2 67.6 67.0
-IN-C (ViT) 49.0 48.8 48.7 48.8
-COMM_INTERVAL 1 10 100 1000
-CIFAR10-C 22.6 22.9 22.6 22.0
-CIFAR100-C 34.7 34.7 34.9 43.2
-IN-C (ResNet) 67.1 67.2 67.2 67.4
-IN-C (ViT) 48.4 48.8 48.8 48.7
-PERT_MAGNITUDE 0.0 0.1 0.2 0.3
-CIFAR10-C 24.3 22.9 22.5 22.2
-CIFAR100-C 36.3 34.7 34.9 34.9
-IN-C (ResNet) 69.3 67.2 67.0 66.9
-IN-C (ViT) 49.6 48.8 48.5 48.6
-VII-C2 Sensitivity Study
+Figure 7: Parameter study with respect to CLUSTER_NUM, KMEANS_SIZE, COMM_INTERVAL & PERT_MAGNITUDE (from top to bottom) under Mixed Distribution Shifts.
+VII-B4 FreDA adapts robustly to hyperparameter variations
 We investigate the impact of the key hyperparameters: CLUSTER_NUM, KMEANS_SIZE, COMM_INTERVAL, and PERT_MAGNITUDE.
-From TABLE VI, we have the following observations: 1) The choice of CLUSTER_NUM influences model performance, especially on more complex datasets.
-While performance remains relatively stable even with a small number of clusters (e.g., two clusters), increasing the number of clusters beyond four tends to lead to slight performance degradation, particularly on datasets like ImageNet-C.
-This suggests that while more clusters can improve adaptation capacity, excessively increasing the number can introduce overfitting and reduce generalization.
+From Fig.
+7, we have the following observations: a) CLUSTER_NUM is largely insensitive in the range of 2–8: performance varies only marginally across CIFAR10/100-C and remains stable on average, with a clear drop emerging only at the extreme setting of 16.
+We attribute the degradation at 16 to over-partitioning, where the model tends to overfit to subtle, cluster-specific variations rather than capturing domain-invariant patterns.
 Therefore, a moderate number of clusters, around 4, appears to strike a good balance between adaptation flexibility and maintaining model robustness.
-2) Varying KMEANS_SIZE from 256 to 2048 results in stable performance across all datasets, indicating that our method is robust to changes in cluster sizes.
-3) Our approach shows general robustness to communication frequency (varying COMM_INTERVAL from 1 to 1,000).
+b) Varying KMEANS_SIZE from 256 to 2048 results in stable performance across all datasets, indicating that our method is robust to changes in cluster sizes.
+c) Our approach shows general robustness to communication frequency (varying COMM_INTERVAL from 10 to 1,000).
 On simpler datasets such as CIFAR10-C, infrequent communication (e.g., interval $f=1000$) performs best, likely due to effective independent learning.
-In contrast, for complex datasets, more frequent communication (e.g., $f=1$) improves performance, likely by mitigating divergence among local branches and ensuring model consistency.
-4) Adjusting the perturbation magnitude (PERT_MAGNITUDE) has a noticeable effect, particularly on simpler datasets.
-Increasing perturbation improves performance, suggesting that spectral augmentation contributes to better generalization.
-However, for more complex datasets such as CIFAR100-C, higher perturbation levels lead to slight decline in performance, likely due to noise disrupting the alignment of finer features.
-A perturbation magnitude of 0.1 appears to strike an effective balance, offering the benefits of augmentation without negatively impacting feature alignment.
-TABLE VII: Classification error rate ($\downarrow$) on CIFAR-10-C (C10), CIFAR-100-C (C100), and ImageNet-C (IN) using ResNet-50 & ViT-Base backbones under Continual Setting, averaged over 15 corruptions.
+In contrast, for complex datasets (e.g.
+CIFAR100-C), more frequent communication (e.g., $f=10$) improves performance, likely by mitigating divergence among local branches and ensuring model consistency.
+d) Introducing perturbation (PERT_MAGNITUDE) from 0 to a positive value yields a clear performance gain, confirming the effectiveness of our frequency-based augmentation strategy.
+Furthermore, when varying the magnitude between 0.1 and 0.3, FreDA maintains robust performance across datasets, indicating its resilience to moderate perturbation levels.
+TABLE V: Classification error rate ($\downarrow$) on CIFAR-10-C (C10), CIFAR-100-C (C100), and ImageNet-C (IN) using ResNet-50 & ViT-Base backbones under Continual Setting, averaged over 15 corruptions.
 Methods C10 C100 IN(ResNet) IN(ViT)
@@
 TBN 20.4 35.4 68.6 -
-TENT (ICLR 21’) 20.0 62.2 62.6 54.5
-ETA (ICML 22’) 17.9 32.2 60.2 49.8
-AdaContrast (CVPR 22’) 18.5 33.5 65.5 57.0
-CoTTA (CVPR 22’) 16.5 32.8 63.1 77.0
-SAR (ICLR 23’) 20.4 32.0 61.9 51.7
-RoTTA (CVPR 23’) 19.3 34.8 67.3 58.3
-RDumb (NeurIPS 23’) 17.8 34.1 90.6 50.2
-DeYO (ICLR 24’) 87.0 98.1 90.6 94.3
-UnMix-TNS (ICLR 24’) 24.9 32.7 75.4 -
+TENT 20.0 62.2 62.6 54.5
+ETA 17.9 32.2 60.2 49.8
+AdaContrast 18.5 33.5 65.5 57.0
+CoTTA 16.5 32.8 63.1 77.0
+SAR 20.4 32.0 61.9 51.7
+RoTTA 19.3 34.8 67.3 58.3
+RDumb 17.8 34.1 90.6 50.2
+DeYO 87.0 98.1 90.6 94.3
+UnMix-TNS 24.9 32.7 75.4 -
 FreDA (ours) 19.5 32.5 60.2 47.9
-VII-C3 Performance under continual settings
+VII-C Performance with Transformer Backbone
+In addition to evaluating our model on commonly compared CNN backbones, we further assess its performance under transformer-based architecture, specifically using the ViT-Base backbone.
+Here, we report results on the ImageNet-C benchmark using ViTBase-LN [49] (see TABLE IV), where the pretrained weights are obtained from torchvision.
+Importantly, all experimental configurations are kept consistent with those used in the CNN-based experiments.
+As shown, our method continues to deliver strong performance, demonstrating its robustness and adaptability across different backbone architectures.
+VII-D Performance under Continual Settings
 Although our method is specifically designed for mixed domain scenarios, we also evaluated its performance under the conventional continual test-time adaptation [5, 3] setting to assess its robustness in different contexts.
@@
 Our results suggest that the proposed FreDA not only excels in its intended mixed domain scenarios but also generalizes effectively to standard continual adaptation tasks, providing a robust solution across various distributional challenges.
-VIII Conclusion and Future Work
-This paper advances Test-Time Adaptation (TTA) by tackling the real-world complexities of heterogeneous data streams.
-Our decentralized approach precisely manages diverse data shifts, improving model adaptation in varied settings.
-By integrating Fourier-based augmentation, we expand the range of confident samples for each distribution shift, further boosting model performance.
-The experimental results underscore the efficacy of FreDA, highlighting its potential to influence the field and guide future research in adapting to dynamic and diverse data shifts.
-While FreDA addresses a critical challenge in handling heterogeneous data streams, providing a solid pipeline for this issue, there are still avenues for further enhancement.
-Our current aggregation approach, which averages models based on cluster counts, has been effective in solving the problem at hand.
-However, exploring alternative strategies – such as weighting models by the divergence between clusters – might lead to incremental improvements.
-Additionally, refining the sample selection process from a original sample-level focus to a more granular patch-level could extend FreDA’s applicability to tasks such as segmentation, further enhancing its versatility in real-world scenarios.
--A Adaptation Scenarios
--A1 Mixed Domains
-: In this scenario, the model processes a long sequence of test samples where each sample $x_{i}\sim p_{e_{i}}(x)$ is independently drawn from a randomly selected target domain $\mathcal{D}_{e_{i}}\in\{\mathcal{D}_{t_{1}},\mathcal{D}_{t_{2}},\dots,\mathcal{D}_{t_{N}}\}$ and a randomly selected class $c_{i}$ among classes $\{1,2,\dots,C\}$.
+VIII More Details and Discussion
+VIII-A Adaptation Scenarios
+VIII-A1 Mixed Domains
+In this scenario, the model processes a long sequence of test samples where each sample $x_{i}\sim p_{e_{i}}(x)$ is independently drawn from a randomly selected target domain $\mathcal{D}_{e_{i}}\in\{\mathcal{D}_{t_{1}},\mathcal{D}_{t_{2}},\dots,\mathcal{D}_{t_{N}}\}$ and a randomly selected class $c_{i}$ among classes $\{1,2,\dots,C\}$.
 The sequence is represented as:
@@
 where each target domain index $e_{i}\in\{1,2,\dots,N\}$ and class number $c_{i}\in\{1,2,\dots,C\}$ are independently and randomly selected for each sample $x_{i}$.
--A2 Continual Domain Adaptation
-: In this setting, the model adapts online to a sequentially presented series of test domains, where each domain shift occurs only once and does not reappear.
+VIII-A2 Continual Domain Adaptation
+In this setting, the model adapts online to a sequentially presented series of test domains, where each domain shift occurs only once and does not reappear.
 The sequence progresses through distinct target domains in a fixed order $D_{1}\to D_{2}\to\dots\to D_{N}$, with samples from each domain appearing contiguously.
@@
 The sequence is structured as:
-$\begin{split}&\underbrace{\{(x_{1}^{D_{1},c_{1}}),(x_{2}^{D_{1},c_{2}}),\dots,(x_{l_{1}}^{D_{1},c_{l_{1}}})\}}_{\text{Samples from domain }D_{1}}\\ &\to\underbrace{\{(x_{l_{1}+1}^{D_{2},c_{l_{1}+1}}),\dots,(x_{l_{1}+l_{2}}^{D_{2},c_{l_{1}+l_{2}}})\}}_{\text{Samples from domain }D_{2}}\\ &\to\dots\to\underbrace{\{(x_{l_{1}+\dots+l_{N-1}+1}^{D_{N},c_{l_{1}+\dots+l_{N-1}+1}}),\dots,(x_{k}^{D_{N},c_{k}})\}}_{\text{Samples from domain }D_{N}}\end{split}$
+[…]
 where:
@@
   • $c_{m}\in\{1,2,\dots,C\}$ is the randomly selected class label for the $m$-th sample, independent of domain transitions.
-Figure 6: Examples from ImageNet-C under common image corruptions. The images showcase a range of corruption types (e.g., noise, blur, and weather distortions) at varying severity levels.
-Figure 7: Samples from DomainNet126 across four subdomains (Real, Sketch, Painting, Clipart). These visualizations reflect the stylistic and perceptual variations inherent in each domain.
--B Dataset Visualization
-To further illustrate the characteristics of the datasets used in our evaluation, we present visualizations of the data distribution across different corruption types (Fig.
-6), natural domain shifts (Fig.
-7), and medical centers (Fig.
-2).
-These figures highlight the diverse challenges that our models face in each evaluation scenario, providing insight into the complexity of the test conditions.
--C Relation to Frequency Domain Learning
+VIII-B Further Remarks on Related Works
+VIII-B1 Relation to Frequency Domain Learning
 Recent advances highlight frequency-based techniques as powerful tools for domain transfer.
@@
 Motivated by these advancements, we discover that frequency information inherently captures domain characteristics, making it a valuable medium for decoupling mixed target domains – an aspect largely unexplored in prior work.
-On this basis, we propose FreDA, the first framework addressing TTA under mixed domain shifts in Fourier space, with a decentralized adaptation and perturbation mechanism.
--D Relation to Multi-Target Domain Adaptation
-TTA under mixed distribution shifts shares similarities with multi-target unsupervised domain adaptation (MT-UDA) [11, 12, 58, 59], yet diverges critically in complexity: while MT-UDA assumes static, predefined target domains and leverages labeled source data for explicit domain alignment, TTA operates with no access to source data and must adapt to dynamic, unpredicTABLE target streams in an online manner.
+On this basis, we propose FreDA, addressing TTA under mixed domain shifts in Fourier space, with a decentralized adaptation and perturbation mechanism.
+VIII-B2 Relation to Multi-Target Domain Adaptation
+TTA under mixed distribution shifts shares similarities with multi-target unsupervised domain adaptation (MT-UDA) [62, 11, 12, 13, 63, 64], yet diverges critically in complexity: while MT-UDA assumes static, predefined target domains and leverages labeled source data for explicit domain alignment, TTA operates with no access to source data and must adapt to dynamic, unpredictable target streams in an online manner.
 This eliminates direct source-target discrepancy computation and demand robust incremental adaptation rather than offline multi-domain optimization.
 These differences highlight the need for novel methodologies tailored specifically to TTA, going beyond the solutions developed for MT-UDA.
--E Relation to Decentralized, Federated, Distributed Learning
+Figure 8: Examples from ImageNet-C under image corruptions. The images showcase a range of corruption types (e.g., noise, blur, and weather distortions) at varying severity levels.
+Figure 9: Samples from DomainNet126 across four subdomains (Real, Sketch, Painting, Clipart). These visualizations reflect the stylistic and perceptual variations inherent in each domain.
+VIII-B3 Relation to Decentralized, Federated, Distributed Learning
 This work also intersects with decentralized, federated, and distributed learning due to splitting data batches into disjoint subsets and applying decentralized model adaptation.
@@
 In our case, as target samples are mixed in a batch, data privacy is not a concern.
-However, similar to federated learning, our approach also involves weight aggregation over subnetworks to enhance their base models before next batch adaptation.
+However, similar to federated learning, our approach also involves weight aggregation over subnetworks to enhance their base models.
 Third, distributed learning aims to improve training efficiency on large-scale datasets by partitioning data for synchronized training [67].
 In contrast, our method operates in a real-time fine-tuning context with limited data at one time, hence scalability is less of a concern.
+VIII-C Dataset Visualization
+To further illustrate the characteristics of the datasets used in our evaluation, we present visualizations of the data distribution across different corruption types (Fig.
+8), natural domain shifts (Fig.
+9), and medical centers (Fig.
+3).
+These figures highlight the diverse challenges that our models face in each evaluation scenario, providing insight into the complexity of the test conditions.
+IX Conclusion
+This paper advances Test-Time Adaptation by tackling the practical challenges posed by heterogeneous data streams.
+We introduce FreDA, a frequency-based decentralized adaptation framework that partitions incoming samples and performs localized adaptation, while employing distribution-aware augmentation to effectively handle mixed distribution shifts.
+Empirical results across multiple benchmarks demonstrate the effectiveness of our approach, highlighting its potential to drive future research on adapting to dynamic and diverse domain shifts.
+Future Work: FreDA provides a principled framework for adapting to mixed distribution shifts, and we believe several promising directions could further enhance its capabilities.
+First, beyond the current proactive clustering, future work could investigate adaptive, learnable partitioning strategies that dynamically refine subset boundaries and better capture subtle relationships among target subdomains.
+Second, our frequency-based augmentation is currently applied at the sample level, which is well-suited for classification.
+Extending it to patch-level operations in spatial–frequency space may enrich the model’s granularity and open new possibilities for fine-grained tasks such as segmentation.
+Together, these directions offer natural extensions to broaden FreDA’s applicability in increasingly realistic and dynamic environments.